### v1

# MM-Interleaved: Interleaved Image-Text Generative Modeling via Multi-modal Feature Synchronizer

Abstract
* Equal contribution.† This work is done when Changyao Tian, Weiyun Wang, and Zhe Chen are interns at Shanghai AI Laboratory.🖂 Corresponding author (daijifeng@tsinghua.edu.cn).
Developing generative models for interleaved image-text data has both research and practical value. It requires models to understand the interleaved sequences and subsequently generate images and text. However, existing attempts are limited by the issue that the fixed number of visual tokens cannot efficiently capture image details, which is particularly problematic in the multi-image scenarios. To address this, this paper presents MM-Interleaved, an end-to-end generative model for interleaved image-text data. It introduces a multi-scale and multi-image feature synchronizer module, allowing direct access to fine-grained image features in the previous context during the generation process. MM-Interleaved is end-to-end pre-trained on both paired and interleaved image-text corpora. It is further enhanced through a supervised fine-tuning phase, wherein the model improves its ability to follow complex multi-modal instructions. Experiments demonstrate the versatility of MM-Interleaved in recognizing visual details following multi-modal instructions and generating consistent images following both textual and visual conditions. Code and models are available at https://github.com/OpenGVLab/MM-Interleaved.
1 Introduction
Interleaved image-text data, where the text and images are related in content and intertwined in layout, is very common on the internet. As an extension of image-text pairs widely used in previous multi-modal models [37, 21, 82, 22, 70], the interleaved format not only covers a broader range of data but also presents more complex information structures. Developing multi-modal models that can simultaneously understand and generate such interleaved image-text data has significant research value and practical application potential. It broadens the application scope of previous multi-modal models [83] and enables more unified processing of multi-modal data, bringing previously isolated research fields such as text-to-image generation [74, 71] and visual question-answering [48, 53] together.
Figure 1: Illustration of multi-modal feature synchronizer. In the auto-regressive generation of interleaved image-text sequences, besides interacting with low-resolution image tokens with self-attention, tokens in MM-Interleaved could also use the multi-modal feature synchronizer to cross-attend multi-scale high-resolution image features. The multi-modal feature synchronizer ensures the cross attention will have exactly the same causal relation between image and text tokens.
With recent progress in multi-modal modeling with Large Language Models (LLMs) [7, 64], exploring how LLMs can be leveraged for interleaved image-text modeling has become increasingly attractive. A core challenge of interleaved multi-modal LLMs is how to effectively handle multiple images within the limited context length of LLMs. In most multi-modal LLMs, images are encoded as visual tokens from an image tokenizer (also known as image encoder) and fed into LLMs together with text tokens [43, 18, 83, 40]. To reduce computational demands and the required context length of LLMs, Perceiver Resamplers [3] are commonly used to map each image from up to 1024 tokens to a fixed number of visual tokens (e.g., 32 or 64 tokens) as in Fig. 2. Due to the relatively small token number, critical image details may be overlooked, especially in tasks requiring fine-grained observation. Increasing the number of visual tokens [55, 99] can address this issue to some extent. While the context length of LLMs is typically limited, allocating a large number (e.g., 2890 in [84]) of visual tokens per image would pose a significant limitation in multi-image scenarios.
We note that the main problem is that the Perceiver Resampler only takes image features as input and is context-insensitive. It compresses each input image to a fixed number of tokens, which makes it harder or even impossible to preserve all information required and fulfill all needs for subsequent generation. In contrast, allowing intermediate layers in decoder networks to dynamically extract a small amount of relevant information from the images can compensate for this information loss. Such a dynamic extraction mechanism should also efficiently handle scenarios requiring multiple images and high-resolution feature maps.
Motivated by this, we designed a fine-grained multi-modal feature synchronizer based on the deformable sparse attention [110]. As shown in Fig. 1, it enables the multi-modal LLMs to observe multi-scale high-resolution image features from multiple images efficiently. This mechanism facilitates the extraction of fine-grained image details with only a small number of input visual tokens, which is particularly applicable to interleaved image-text scenarios with multiple images.
Based on this module, we propose MM-Interleaved, a novel end-to-end generation model for processing interleaved image-text data, as shown in Fig. 3. Images and text are first mapped into tokens through their respective tokenizers (i.e., image encoder and word embedding) and then fed to the LLM, arranged in their original order. A special token <BoI> is introduced to represent “Begin of Image”. When the LLM processes image and text sequences, each token in the intermediate layers directly observes multi-image and multi-scale feature maps from the previous context through the proposed feature synchronizer module. After being processed by the LLM, the features of the text tokens predict the next text word. When the <BoI> token is predicted, an image decoder based on diffusion models is used to predict the next image. All previous LLM output features are passed into the image decoder to generate images. Each pixel in the intermediate layers of the decoder can also extract details of the previous image through the proposed synchronizer module.
Our models are pre-trained on a mixture of image-text pairs and interleaved image-text sequences. We do not use any in-house data. Similar to previous multi-modal LLMs, supervised fine-tuning can further enhance our model capabilities. Thanks to end-to-end generative modeling, fine-tuning can be applied to both text and image generation. Our model is fine-tuned on several tasks, including visual question-answering, image caption, referring expression grounding, text-to-image generation, segmentation-to-image translation, and visual storytelling. Compared with previous image-to-text and text-to-image methods, our model achieves competitive results and shows higher token efficiency. Compared with joint image and text generation models, we set the new SotA results for a wide range of tasks. Overall, MM-Interleaved can handle interleaved image-text inputs and recognize fine-grained details across multiple images to produce accurate textual descriptions and visually consistent images, demonstrating the potential of high-performance multi-modal models that can understand and generate interleaved image-text content.
Figure 2: Comparison of different types of image-text generative modeling. (a) and (b) can only generate text. (c) and (d) can generate both images and text. All types except (d) are limited by the fixed number of visual tokens extracted by context-insensitive Resampler, which can not efficiently capture image details and is problematic in multi-image scenarios.
2 Related Work
Modeling of Paired Image-Text Data plays an important role in the progress of multi-modal research in recent years. [37, 24, 76] has released a series of public large-scale image-text pairs datasets. Based on these data, models trained by image-text contrastive learning [70, 39, 82, 50] have been proven able to recognize and understand open-world semantics. Subsequent work [93, 111, 46, 97] further incorporates the text generation tasks such as image captioning. Some other works [74, 71] were proposed to generate images based on text prompts. The latest progress of LLMs [64, 65] has created a new era, leading to the emergence of many LLM-based multi-modal models trained using image-text pairs [48, 53, 90, 91]. For example, [53] connect pre-trained LLMs and vision foundation models with linear projections to build powerful multi-modal text-generation models.
Figure 3: Architecture of the proposed MM-Interleaved Model. The red lines represent how the multi-scale image features are generated and utilized. MM-Interleaved incorporates image encoder to both extract high-resolution multi-scale image features and map each image into a fixed number of low-resolution visual tokens. These visual tokens are fed into the multi-modal LLM along with text tokens. LLM uses the feature synchronization module to extract high resolution image details, and auto-regressively generates text tokens. After that, a DM-based image decoder generates next image conditioned on the previous context features from LLM as well as the multi-scale image features from the feature synchronization module.
Modeling of Interleaved Image-Text Data has received increasing attention recently. [33, 3] were the first to focus on understanding such data relying on non-public datasets. To promote the development of this field, [109, 45] release public and large-scale interleaved image-text datasets. More works [103, 101] have begun to focus on understanding the interleaved data. Nevertheless, their generative capabilities remain limited to text. [43, 99, 83, 18] have initially attempted to generate images and text in the interleaved context concurrently. [43, 83] involve a two-stage generation process, where text and image are only generated in the first and second stages, respectively. [99] uses VQ-VAE [86] to convert images into discrete tokens, facilitating token-level auto-regressive modeling as traditional language modeling. However, it primarily focuses on image generation capabilities and exhibits notable weaknesses in image understanding. DreamLLM [18], a concurrent work to ours, also focuses on single-stage end-to-end modeling using raw image pixels as input. Despite these efforts, they only feed image information at the input of LLMs, which are limited by the problem that fixed number of visual tokens cannot efficiently describe image details.
Integrating Image Details into LLMs is important for multi-modal LLMs. Most works use Perceiver Resamplers [3, 48, 107] to extract image information via cross-attention, mapping each image into a fixed number of visual tokens. For example, [3] equips Resamplers in the intermediate layers of LLMs, injecting extracted image features into LLMs through gated residual connections. [48, 107] introduces Resamplers at the bottom of LLM to insert the extracted visual tokens into the input text sequences. While these methods achieve good performance, image details may be overlooked due to the small number (e.g., 32 or 64) of visual tokens. To preserve more image details, [6, 55, 53, 10] increase the number of input visual tokens per image to hundreds. [84] further increases this number to 2,890. Although it mitigates information loss, the computational and memory demands of LLMs are significantly increased. Increasing the number of visual tokens per image is particularly problematic in multi-image scenarios, where multiple images naturally require more visual tokens, making it hard to use for interleaved image-text data.
3 Method
3.1 Task Formulation
To build an end-to-end generative model for interleaved image-text data. We first consider an interleaved image-text sequence $X=\{x_{1},x_{2},x_{3},\dots\}$, where each element $x_{n}$ is either a text token (denoted as $x_{n}^{L}$) or a whole image (denoted as $x_{n}^{V}$). Text and images are arranged in the order in which they appear in the original content. In multi-modal LLMs, a common practice is first extracting embedding for each text token and each image before being fed into LLMs, i.e., $e_{n}^{L}=\mathcal{E}_{L}(x_{n}^{L})$ and $e_{n}^{V}=\mathcal{E}_{V}(x_{n}^{V})$, where $\mathcal{E}_{L}$ denotes word embedding following standard practices in NLP. $\mathcal{E}_{V}$ is typically an image encoder (e.g., ViTs [19]) followed by a Perceiver Resampler [3] to map each image to a fixed number of tokens. Then, the generative modeling is trained to maximize the log-likelihood of the multi-modal sequence, given by:
$\log p(X)$ $=\sum_{n}\log p(x_{n}|x_{1},x_{2},\dots,x_{n-1})$ (1)
$=\sum_{n}\log p(x_{n}|e_{1},e_{2},\dots,e_{n-1})$
$=\sum_{n\in\mathcal{I}_{L}}\log p(x_{n}^{L}|e_{<n})+\sum_{n\in\mathcal{I}_{V}}\log p(x_{n}^{V}|e_{<n}),$
where $\mathcal{I}_{L}$ and $\mathcal{I}_{V}$ represent the index sets for text tokens and images within the sequence, respectively. $<n$ in the subscript represents the abbreviation of $\{1,2,\dots,n-1\}$. The following paragraphs provide explanations of Eq. (1).
Text Generation with Multi-modal Condition. The log-likelihood $\log p(x_{n}^{L}|e_{<n})$ is similar to traditional causal language modeling, except that the condition also includes previous images. Recent works [48, 53, 90, 91] have demonstrated the effectiveness of using LLMs for processing additional visual inputs. They apply a linear classifier on top of the LLMs. The loss function for text generation is
$L_{\text{NTP}}(x_{n}^{L}|e_{<n})=-\bar{x}_{n}^{L}\cdot\log\mathrm{softmax}\big{(}W\cdot\mathcal{D}_{\text{LM}}(e_{<n})\big{)},$ (2)
where $W$ is the linear classification weight, $\mathcal{D}_{\text{LM}}$ denotes the LLM network (e.g., LLaMA [85]), $\bar{x}_{n}^{L}$ is the one-hot vector indicating the ground-truth text token.
Image Generation with Multi-modal Condition. Maximizing the log-likelihood $\log p(x_{n}^{V}|e_{<n})$ aligns with the denoising-diffusion process [54], which recently achieves widespread success in image generation. When conditional inputs $x_{<n}$ are all text, this is identical to text-to-image diffusion models, where maximizing the log-likelihood is derived as minimizing the diffusion modeling loss as
$L_{\text{NIP}}(x_{n}^{V}|e_{<n})=\mathbb{E}_{\epsilon,t}~{}||\epsilon-\mathcal{D}_{\text{DM}}\big{(}x_{n,t}^{V},t,\mathcal{D}_{\text{LM}}(e_{<n})\big{)}||^{2},$ (3)
where $\mathcal{D}_{LM}$ is the language model encoding text inputs, $\mathcal{D}_{DM}$ is the diffusion model for image denoisiong. $x_{n,t}^{V}$ is the noisy version of the original image at the denoising step $t$, and the denoising network $\mathcal{D}_{DM}$ is trained to predict the noise $\epsilon$. We found that such modeling is also applicable when conditional input is interleaved image-text data.
Note that our modeling allows for the flexible combination of different language models, image encoders, and image decoders to fully leverage a variety of pre-trained models. It does not need to insert additional special tokens specifically for image generation into LLMs. The whole framework can be optimized end-to-end.
3.2 Architecture
Building upon the task formulation in Sec. 3.1, we propose a novel multi-modal architecture for processing interleaved image-text data. It integrates a Visual Foundation Model (VFM), a Large Language Model (LLM), and a Diffusion Model (DM). This integration aims to excel in both understanding and generation tasks of text and images by leveraging the strengths of each model type. As illustrated in Fig. 3, our architecture comprises three key components:
(1) VFM-based Image Tokenizer $\mathcal{E}_{V}$ that maps each image $x^{V}\in\mathbb{R}^{H\times W\times 3}$ (e.g., $H=W=224$) into a fixed number of visual tokens $e^{V}\in\mathbb{R}^{N\times C}$ ($N=32$ by default). $C$ is the channel dimension. It consists of a pre-trained vision foundation model (e.g., CLIP-ViT [70]) for feature extraction and a Perceiver Resampler [3] to reduce the number of visual tokens. We also use a ViT-Adapter [12] to extract multi-scale image features $F^{V}\in\mathbb{R}^{(\sum_{i=1}^{L}H_{i}\times W_{i})\times C}$ for fine-grained feature fusion in subsequent networks, where $L=4$ and $H_{i}=\frac{H}{2^{i+1}},W_{i}=\frac{W}{2^{i+1}}$ by default.
(2) LLM-based Multi-modal Model $\mathcal{D}_{\text{LM}}$ that extracts context features from the interleaved image-text sequences. A pre-trained LLM (e.g., Vicuna [105]) is utilized. Its input sequence $E\in\mathbb{R}^{K\times C}$ is a concanation of embeddings $(e_{1},e_{2},\dots)$, where $e_{n}$ is either a word embedding $e_{n}^{L}\in\mathbb{R}^{1\times C}$ or an image embedding $e_{n}^{V}\in\mathbb{R}^{N\times C}$. $K$ is the total number of input tokens. We also introduce a fine-grained feature synchronizer module for letting the intermediate layers in $\mathcal{D}_{\text{LM}}$ directly access and extract multi-scale image features on demand.
(3) DM-based Image Decoder $\mathcal{D}_{\text{DM}}$ that generates the image based on previous image-text sequences. A pre-trained diffusion model (e.g., Stable Diffusion [74]) is utilized. To provide the conditional inputs for $\mathcal{D}_{\text{DM}}$, Resampler [3] is employed to map the output features from LLM to a fixed number of conditional tokens. The fine-grained feature synchronizer module is also used here for providing detailed visual conditions, which is very useful for tasks requiring visual alignment (e.g., image translation).
Then we introduce the details of the proposed architecture.
Figure 4: Architecture of the MMFS module. The query feature is passed through a linear layer and added to the image index embedding. Two linear layers are used to predict the sampling offsets and unnormalized attention weights of each image respectively. The sampling offsets are added with the query’s reference point to form the corresponding sampling locations, which are shared and rescaled across all feature maps of the same image. The output feature is the weighted sum of the sampled spatial features from these points.
Multi-image and Multi-scale Feature Synchronizer (MMFS). The MMFS module is designed to dynamically capture image details from multiple images in an interleaved context. It leverages the Deformable Attention [110] mechanism to achieve efficient and sparse image attention. As is shown in Fig. 4, given a query token that requires detailed image features, the MMFS module only attends to a small set of sampling points around a reference point on the reference images. Let $f_{q}\in\mathbb{R}^{C}$ represent the features of the query token. $\hat{p}_{q}\in[0,1]^{2}$ denotes the relative image coordinate of its reference point. By default, $\hat{p}_{q}=(0.5,0.5)$ corresponds to the center of the images. $\{F^{V}_{m}\}_{m=1}^{M}$ are the multi-image multi-scale feature maps, where $F^{V}$ is the multi-scale feature extracted by the image tokenizer, $M$ is the number of reference images. The output feature $f_{o}\in\mathbb{R}^{C}$ is given by
$q^{(m)}$ $=W_{q}\cdot f_{q}+\text{PosEmbed}(m),$ (4)
$p_{q}^{(m)}$ $=W_{p}\cdot q^{(m)}+\hat{p}_{q},\ \ A_{q}^{(m)}=W_{A}\cdot q^{(m)},$
$p_{q}$ $=\mathrm{Concat}(p_{q}^{(1)},\cdots,p_{q}^{(M)}).$
$A_{q}$ $=\mathrm{softmax}\left(\mathrm{Concat}(A_{q}^{(1)},\cdots,A_{q}^{(M)})\right).$
$f_{o}$ $=\mathrm{DeformAttn}(\{F^{V}\}_{m=1}^{M},A_{q},p_{q}),$
where $\text{PosEmbed}\in\mathbb{R}^{\bar{M}\times C}$ is a learnable positional embedding, $m$ indexes from the maximum of $\bar{M}$ reference images. $W_{q},W_{p},W_{A}$ are learnable linear projection weights. The coordinate of sampling points $p_{q}^{(m)}$ and the corresponding attention weights $A_{q}^{(m)}$ are first calculated for each image separately. Then, the attention weights are normalized among different images via the softmax function. The $\mathrm{DeformAttn}$ operator extracts feature at coordinates $p_{q}\in\mathbb{R}^{M\times L\times K\times 2}$ from the corresponding feature map, and performs a weighted summation according to $A_{q}\in\mathbb{R}^{M\times L\times K}$. Here, $L$ and $K$ denote the number of multi-scale feature levels and sampling points per feature map, respectively.
MMFS can be applied to both image and text decoding, avoiding information bottlenecks caused by the widely used Resamplers in multi-modal LLMs. It is especially efficient for processing multiple high-resolution images in context.
Multi-modal LLM with MMFS. The input of multi-modal LLMs is an interleaved sequence of image and text token embeddings, which always starts with a special token <s> and ends with another special token </s>. Image token embeddings are inserted at the corresponding positions in the original sequence. Another special token <BoI> is added in front of each image to represent “Begin of Image”.
MMFS modules are inserted between the self-attention layer and the feed-forward layer of the LLM every fixed number of blocks. Query token $f_{q}$ iterates over each token in the LLM, which can only access the previous images. $\hat{p}_{q}=(0.5,0.5)$. The output of MMFS is multiplied by tanh($\alpha$) before being added back to $f_{q}$, where $\alpha$ is a zero-initialized learnable scalar.
Image Decoder with MMFS. The output features from the multi-modal LLM are further processed by another Perceiver Resampler before being fed into the image decoder. For each image to be generated, the Resampler maps previous context features to a fixed number of tokens (e.g., 77 tokens) to match the pre-trained diffusion model.
MMFS modules are inserted after each upsampling block in the U-Net structure of the diffusion model. Query token $f_{q}$ iterates over each pixel in the feature map. $\hat{p}_{q}$ is set as the spatial coordinate of the query pixel. The output of MMFS is further processed by a zero-initialized convolution layer before being added back to $f_{q}$.
Training Target and Inference Pipeline. The training objective is defined as the sum of Next-Text-Token Prediction loss in Eq. (2) and Next-Image Prediction loss in Eq. (3),
$\mathcal{L}=\mathcal{L}_{NTP}+\lambda~{}\mathcal{L}_{NIP}$ (5)
where $\lambda$ is a hyperparameter used to determine the relative loss weight between the image and text decoding branches. The whole framework can be optimized end-to-end.
During inference, the images and texts are generated in an auto-regressive manner. Text tokens are sampled from the distribution predicted by the multi-modal LLM. When the generated token is <BoI>, the diffusion model is called for generating the next image.
4 Experiment
4.1 Implementation Details
[…]
Models for Text-Generation Only
MetaLM [30] MetaLM - - - 82.2 43.3 58.7 41.1 11.4 41.1 11.4
OF-9B [5] MPT-7B - - - 79.5 59.5 52.7 37.8 27.5 24.2
IDEFICS-80B [36] LLaMA-65B - - - 91.8 53.7 65.0 60.0 45.2 36.0 30.9
KOSMOS-1 [33] MetaLM H - - 65.2 46.7
KOSMOS-2 [67] KOSMOS-1 H - - 66.7 45.6
Flamingo-9B [3] Chinchilla-7B H - - 79.4 61.5 51.8 44.7 28.8 31.8 48.0
Flamingo-80B [3] Chinchilla-70B H - - 84.3 67.2 56.3 50.6 31.6 35.0 52.0
IDEFICS-80B-I [36] LLaMA-65B - I - 117.2 65.3 104.5 37.4 26.0
mPLUG-DocOwl [95] LLaMA-7B - I A 52.6 62.2 57.4
BLIP-2 [48] Vicuna-7B - I A 74.9 107.5 38.6 25.3 40.1
BLIP-2 [48] Vicuna-13B - I A 71.6 103.9 41.0 41.0 19.6 42.5
InstructBLIP [14] Vicuna-7B - I A 82.4 123.1 49.2 34.5 50.1
InstructBLIP [14] Vicuna-13B - I A 82.8 121.9 49.5 33.4 50.7
Shikra [10] Vicuna-13B - I A 117.5 73.9 77.4
LLaVA-1.5 [52] Vicuna-7B - I A 78.5 62.0 50.0 58.2
LLaVA-1.5 [52] Vicuna-13B - I A 80.0 63.3 53.6 61.3
VL-GPT-I [108] LLaMA-7B - I A 133.7 67.2 50.3 51.5 38.9 51.8
Qwen-VL [6] Qwen-7B H I A 85.8 121.4 78.8 59.3 35.2 63.8
Qwen-VL-Chat [6] Qwen-7B H I A 81.0 120.2 78.2 57.5 38.9 61.5
Models for both Image and Text Generation
CM3Leon [99] H - - 61.6 10.5 47.6 23.8 37.6 22.6
Emu [83] Vicuna-13B H - - 112.4 52.0 38.2 34.2 47.4
Emu-I [83] Vicuna-13B H - - 117.7 40.0 34.7 35.4 48.0
Emu2 [81] LLaMA-33B H - - 33.3 26.7 40.4 26.2
DreamLLM [18] Vicuna-7B - I - 115.4 17.4 56.6 44.3 38.1 34.9
VL-GPT [108] LLaMA-7B - - - 116.4 51.7 35.8 34.6 34.7 49.9
MM-Interleaved Vicuna-13B - - - 129.0 85.8 106.4 23.5 57.0 40.0 40.8 37.2 48.7
MM-Interleaved-SFT Vicuna-13B - I A 140.5 93.0 123.2 30.3 80.2 51.7 60.5 54.9 61.0 53.7
[…]
Network. We adopt CLIP-ViT-L/14 [70], Vicuna-13B [105] and Stable Diffusion v2.1 [74] as the image encoder, large language model, and image decoder, respectively. For the multi-modal LLM, a Perceiver Resampler with 12 blocks is used to reduce the number of visual tokens per image to 64. The MMFS module is inserted every 4 blocks in the LLM. For the image decoder, a Perceiver Resampler with only 1 block is used to reduce the number of conditional tokens per image to 77. The MMFS module is inserted after each up-sampling block in the image decoder.
Pre-training. Our models are pre-trained on a mixture of interleaved image-text sequences and image-text pairs, including MMC4 [109], LAION-2B [76], LAION-COCO [77], CC-12M [8] and Objects365 [78]. For CC-12M [8] and Objects365 [78], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [48] to caption the images. The sampling probability of MMC4 is twice that of other image-text pair datasets. No in-house data is used.
The images are inserted before or after the corresponding text sentence with equal probability. To optimize training efficiency and data utility, multiple image-text pairs or interleaved image-text sequences are concatenated into extended sequences with the maximum context length (i.e., 2,048 tokens).
The model is pre-trained for 15,000 steps with 11 billion tokens visited. The image encoder and LLM are frozen. The learning rate is set to be $10^{-5}$ for the image decoder and $10^{-4}$ for the rest trainable parameters. The input and output image resolutions are $224\times 224$ and $512\times 512$, respectively.
Supervised Fine-tuning. After pre-training, like multi-modal LLMs, our model capabilities can be further enhanced using supervised fine-tuning. The model is fine-tuned on four types of downstream tasks, which are 1) visual question-answering and image caption, 2) referring expression comprehension, 3) segmentation-to-image translation, and 4) visual storytelling.
More details of datasets and hyper-parameters used for pre-training and fine-tuning could be found in the Appendix.
Model RefCOCO [42] RefCOCO+ [58] RefCOCOg [58]
Val Test-A Test-B Val Test-A Test-B Val Test
OFA-L [88] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.50
VisionLLM-H [90] - 86.70 - - - - - -
Shikra [10] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
MiniGPT-V2 [9] 88.69 91.65 85.33 79.97 85.12 74.45 84.44 84.66
Ferret [96] 89.48 92.41 84.36 82.81 88.14 75.17 85.83 86.34
 Qwen-VL [6] 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
MM-Interleaved (ours) 89.92 92.59 86.54 82.99 88.57 77.07 85.21 84.92
Table 2: Supervised fine-tuning results on referring expression comprehension task. “*” denotes using an additional self-constructed grounding dataset and trained with an image resolution larger than 224.
Model MS-COCO LN-COCO
Text-to-Image Specialists
Retrieval Result 17.97 33.59
DALL-E [71] $\sim$28 -
CogView [16] 27.10 -
CogView2 [17] 24.00 -
Stable Diffusion [74] 12.43 34.26
GLIDE [63] 12.24 -
Make-A-Scene [25] 11.84 -
DALL-E 2 [72] 10.39 -
Muse-3B [94] 7.88 -
Imagen-3.4B [75] 7.27 -
Parti-20B [98] 7.23 15.97
Models for both Image and Text Generation
CM3-13B [1] 29.56 -
VL-GPT [108] 12.25 -
GILL [43] 12.20 -
Emu-13B [83] 11.66 -
CM3Leon-7B [99] 10.82 -
DreamLLM-7B-Stage1 [18] 8.76 22.42
DreamLLM-7B [18] 8.46 20.53
MM-Interleaved-13B (ours) 7.90 23.88
Table 3: Zero-shot text-to-image generation results on MS-COCO [51] and LN-COCO [69]. FID [31] is reported.
4.2 Evaluation
We evaluate the zero-shot visual-language capability of the pre-trained MM-Interleaved as well as its performance on various downstream tasks after fine-tuning and report the results in Tab. 1.
Zero-shot Results after Pre-training. Our method demonstrates strong multi-modal zero-shot comprehension on various benchmarks, including image captioning on COCO [11], Flicker30K [68], NoCaps [2] and Image2Paragraph [44]; general visual question answering on VQAv2 [27], OKVQA [59], VizWiz [29] and TextVQA [80]. We also evaluate on the VisDial [15] visual dialogue task. In the fully decontaminated setting where the image and text in downstream tasks are unseen during pre-training, our model significantly outperforms MetaLM [30], OF-9B [5] and IDEFICS-80B [45] on all captioning, VQA, and multi-modal tasks. This demonstrates the effectiveness of the proposed MM-Interleaved approach. Furthermore, our model even exceeds most trained on vast in-house data like Flamingo-9B [3] and Emu [83], showing the importance of proper architecture for image-text interaction.
Additionally, we evaluate text-conditional image generation on MS-COCO [51] and LN-COCO [69]. On MS-COCO, we sample 8 images per text condition and use CLIP [70] to rerank based on text-image similarity. CLIP reranking is not used for LN-COCO. FID [31] is used to evaluate both datasets. As show in Tab. 3, our model shows competitive text-to-image generation compared to existing image and text generation models.
Fine-tuning Results of VQA and Captioning. Our fine-tuned model achieves state-of-the-art performance on image captioning benchmarks. On visual question answering tasks, it matches the previous best LLaVA-1.5 model [52]. Compared to LLaVA-1.5, our model has two key advantages 1) our method is capable of generating both images and text, while LLaVA-1.5 can only generate text; 2) LLaVA-1.5 uses 576 visual tokens as LM inputs, whereas we only require 64. Our model achieves competitive image understanding with far fewer visual tokens, making it better suited for multi-image scenarios.
Fine-tuning Results of Referring Expression Comprehension. Our model demonstrates state-of-the-art performance on various referring expression comprehension (REC) benchmarks, as shown in Tab. 2. Compared to recent REC models like VisionLLM [90] and Shikra [10], our model outperforms on most datasets. Remarkably, even though we only use public REC data [42, 58] for fine-tuning, our model matches QWen-VL [6] which utilizes a extra 22M in-house grounding dataset and trains at higher resolution (448 vs 224 pixels). This shows the effectiveness of synchronized fine-grained image features for enhancing the REC capability of our model.
Fine-tuning Results of Segmentation-to-Image Translation. We follow the protocol proposed in ControlNet [100] by generating images from ADE20K ground truth semantic masks and use OneFormer [38] for segmenting the generated images. Standard mIoU is reported as a metric for measuring the semantic and structural fidelity of the generated images. We show the results in Tab. 3(a). MM-Interleaved clearly outperforms other baselines including ControlNet by a large margin. Compared to ControlNet, our model could leverage the better representations learned by our generative modeling from large-scale pre-training data, which benefits the image generation tasks. Moreover, thanks to MMFS, our model is capable of generating realistic images with pixel-level precise alignment from a semantic label map, where simply relying on coarse Perceiver Resampler features fails on this task, which we will show in the ablation studies.
ADE20K (GT) [38] VQGAN [20] LDM [73]
0.58 0.21 0.31
PIPT [89] ControlNet [100] Ours
0.26 0.35 0.44
(a) Segmentation-to-image generation results on ADE20K [106] mIoU ($\uparrow$) is reported.
[…]
GILL [43] 0.64 -
MiniGPT-5 [104] 0.70 59.5
MM-Interleaved 0.70 39.7
(b) Image generation results on VIST [34].
Model Pororo Flintstones
StoryDALL-E [57] 25.9 26.5
AR-LDM [66] 17.4 19.3
ACM-VSG [23] 15.4 18.4
MM-Interleaved 14.7 18.7
(c) Autoregressive image generation results on Pororo [49] and FlintStones [28]. FIDs($\downarrow$) are reported.
Table 4: Comparative analysis of image generation ability of MM-Interleaved.
Fine-tuning Results of Visual Storytelling. Besides text-to-image and image-to-image generation, our MM-Interleaved could also condition on interleaved image-text as context for generating new images. We evaluate our model’s performance both for generating the last image on the Visual StoryTelling (VIST) dataset [34] and for generating multiple images autoregressively on the Pororo [49] and Flintstones [28] datasets. As is shown in Tab. 3(b) and  3(c), our model achieves SoTA performance on both CLIP-Similarity and FID metrics compared to previous works on VIST and achieves competitive performance compared with previous specialist methods on Pororo and Flinestones. More qualitative results and implementation details can be found in the appendix.
[…]
32 107.0 32.2 28.7 22.5
32 110.6 30.0 29.8 27.7
256 110.7 32.7 29.2 23.6
256 110.9 29.5 29.6 27.8
(a) pre-training with 224 input resolution
[…]
110.5 30.3 29.9 24.9
115.2 30.5 30.6 30.8
(b) fine-tuning with 448 input resolution
w/ MMFS ADE20$\uparrow$
5.3
35.9
(c) fine-tuning for image translation
Table 5: Ablation on the usage of MMFS module. “Genetation” is the text-to-image generation task. “ADE20k” is the segmentation-to-image translation task. Others are text generation tasks. “# Token” indicates the number of input visual tokens for LLM (32 by default).
[…]
Deformable None $16\times 16$ 110.6 30.0 29.8 27.7
Dense None $16\times 16$ 108.5 30.6 28.4 23.6
Dense Resampler 32 tokens 107.2 30.7 28.9 24.0
Table 6: Ablation on the design choice of MMFS module. Different attention modules can be used in MMFS. We also ablate whether to add additional transition layer before feeding image features into MMFS. Transformer attention with Resampler transition on single scale and single image is similar to the cross attention used in Flamingo. “Self-Attention” has the same parameters with “Resampler”.
4.3 Ablation Study
In the ablation study, we use CLIP-ViT-L/14 [70] as the vision foundation model, OpenLLaMA-3B v2 [26] as the multi-modal LLM and miniSD [^1] https://huggingface.co/justinpinkney/miniSD as the image decoder. The output resolution of each image is 256.
For simplicity, in this subsection, we study the effect of token efficiency, image resolution, cross-attention mechanisms, and the role of MMFS in image generation in a single-image setting, where each token only attends to the nearest preceding image. Please refer to the appendix for more detailed ablations studies in the multi-image setting.
The model is pre-trained for 10k steps on 80 NVIDIA 80G-A100 GPUs. The mini batch size of LAION-COCO [77], LAION-2B [76], and MMC4 [109] on a single GPU is 2, 2, and 4 respectively. For token efficiency, image resolution, and cross-attention mechanism, we evaluate the model’s zero-shot performance on three representative tasks and four datasets, i.e., image caption on the COCO Karpathy test set, text-to-image generation on the COCO Karpathy test set, visual question answering on OKVQA [59], and TextVQA validation set. We use CIDEr, FID-5k, and top-1 accuracy as the evaluation metrics. For cross-attention in image generation, we perform supervised fine-tuning on the ADE20K [106] for the label-to-image generation task.
Token Efficiency. As shown in Tab. 4(a), our method achieves better overall performance while only using 32 visual tokens for each input image. Such results demonstrate the effectiveness of our method when the context length is limited. To avoid the potential information bottleneck caused by the Resampler, we further increase the number of visual tokens to 256. Our method still outperforms the baseline in this setting.
Scaling up Input Image Resolutions. The performance gap between our method and the baseline becomes larger when further increasing the input image resolution from 224 to 448, as shown in Tab. 4(b). Such results indicate our method could better exploit the additional information gained from high resolution.
Comparison between Different Cross Attention Mechanisms. The ablation of adopting different cross-attention mechanisms for LLM is shown in Tab. 6. The overall performance drops when directly replacing MMFS with the vanilla dense cross attention, possibly due to its slower convergence speed on 2D data, and using a resampler performs equally well. We highlight that the model with deformable attention performs significantly better than other attention mechanisms on TextVQA, indicating that deformable attention could effectively and efficiently capture fine-grained information like text needed for the task, such as visual question answering in this case.
MMFS for Image Generation. We study whether adding MMFS is critical for image generation on the segmentation-to-image task. This task is hard as it requires precious pixel-level information to align the given segmentation condition and image output properly. From the results in Tab. 4(c), it is clear that the model without MMFS failed to generate images spatially aligned with the segmentation image condition, showing extremely low mIoU results. We show visualizations of the generated results in the supplementary material to provide a visual comparison. It is worth noting that the quality of the generated image from the model without MMFS is still high and properly aligned with the text description input; however, without MMFS, the model cannot preserve all spatial information, and the spatial alignment for the generated results is poor.
5 Conclusion
In conclusion, this paper presented MM-Interleaved, an end-to-end generative model specifically designed for interleaved image-text data. The unique contribution of MM-Interleaved lies in the novel multi-scale and multi-image feature synchronizer module, which significantly enhances the model’s capability to access and integrate fine-grained image features. This design addresses the longstanding limitation in existing models where a fixed number of visual tokens inadequately capture intricate image details, especially in scenarios involving multiple images. Pre-trained on diverse image-text datasets and further refined through supervised fine-tuning, MM-Interleaved shows a respectful proficiency in processing complex multi-modal instructions. The empirical results modestly highlight its ability to discern intricate visual details and generate images that align accurately with both textual and visual inputs.
Appendix A Additional Ablation Studies
This section provides more ablation studies for MM-Interleaved, all of which share the same settings as those in the main text by default. Note that these models are not fine-tuned on downstream tasks.
Pre-training with Different Loss Terms. As showed in Tab. 7, MM-Interleaved achieves better performance when jointly trained with both $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ on comprehension and generation tasks, demonstrating the mutual benefits between the two loss terms. Moreover, setting $\lambda=10$ achieves a better balance between $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ empirically.
[…]
$\mathcal{L}_{NTP}+100~{}\mathcal{L}_{NIP}$ 106.2 31.1 29.8 24.5
$\mathcal{L}_{NTP}+10~{}\mathcal{L}_{NIP}$ 110.6 30.0 29.8 27.7
$\mathcal{L}_{NTP}+\mathcal{L}_{NIP}$ 110.0 31.4 29.3 26.0
$\mathcal{L}_{NTP}$ only 105.7 29.9 27.6
$\mathcal{L}_{NIP}$ only 34.2
Table 7: Pre-training with different loss terms.
The Relationship between MMFS and resampler. The results in Tab. 8 validate the mutual benefits between our proposed MMFS module and the Resampler used in image tokenizer, as removing either of them leads to an overall performance degradation.
[…]
110.6 30.0 29.8 27.7
107.0 32.2 28.7 22.5
102.7 32.0 27.3 22.0
Table 8: The complementary relationship between MMFS and Resampler. When not using Resampler, we directly feed 32 randomly-initialized learnable embeddings as input visual tokens into the LLM.
MMFS with Multi-Image and Multi-Scale. As shown in Tab. 9, adding multi-image and multi-scale for MMFS improves the model’s performance on VQA tasks. To better take advantage of the multi-image mechanism, we follow [3, 83] to further evaluate our model with few-shot prompts. Two training samples including both image and text are selected as the input prompts. The results in Tab. 10 demonstrates that MMFS with multi-image and multi-scale benefits the most from the few-shot setting, indicating its potential capability for interleaved image-text modeling.
[…]
110.6 30.0 29.8 27.7
111.2 29.5 30.3 28.1
111.2 29.9 31.1 28.2
finetuning with 448 input resolution
115.2 30.5 30.6 30.8
115.4 30.1 31.0 31.3
115.8 30.0 31.7 32.0
Table 9: Zero-shot evaluation results of MMFS.
[…]
36.5 30.7
38.5 31.1
38.7 33.1
finetuning with 448 input resolution
38.3 36.6
38.7 37.7
39.1 37.9
Table 10: 2-shot evaluation results of MMFS on VQA tasks.
Appendix B Implementation Details
B.1 Pre-training
Dataset Details. We use MMC4 [109], LAION-2B, LAION-COCO [77], CC-12M [8] and Objects365 [78] as the pre-training dataset. LAION-2B is the English subset of LAION-5B [76], and we further filter it based on additional metrics including aesthetics scores. LAION-COCO  [77] is a 600M subset of LAION-2B, which is captioned by pre-trained BLIP [47] models. Text prompts with length shorter than 10 are also filtered out. For CC-12M [8] and Objects365 [78], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [48] to caption the images. Following previous works [18, 83], additional filtering rules are appled to the MMC4 dataset [109]. Specifically, images with a CLIP similarity score below 0.24 will be discarded, and only 6 images at most will be kept for each document. We also exclude 100% of all documents that do not contain any images, and 50% of documents that contain only 1 image.
Data Concatenation Strategy. For image-text-pair datasets (e.g., LAION-2B, LAION-COCO [77]), we randomly sample multiple image-text pairs from the same dataset and concatenate them to the maximum context length (i.e., 2048) during pre-training. For interleaved image and text datasets (i.e., MMC4 [109]), we also split and concatenate the documents to form the training samples. Such concatenation strategy can utilize the full context window of Large Language Models and thus achieve high data efficiency.
More Training Details . The detailed hyper-parameters of pre-training are listed in Tab. 11. Besides that, for image generation, we ignore the training loss of images which are the first element in the sequence. The text condition of the rest images are dropped with a 10% probability to improve classifier-free guidance sampling [32].
Hyper-parameters Value
Input image resolution $224\times 224$
Output image resolution $512\times 512$
VFM CLIP-ViT-L/14 (frozen)
LLM Vicuna-13B v1.3 (frozen)
DM Stable Diffusion v2.1
$\lambda$ 10
Cross-attention frequency 4
LLM Multi-scale feature maps $i=2,3,4$
Learning rate 1e-5 (image decoder) 1e-4 (others)
Weight decay 0.05
Warmup steps 1k
Learning rate schedule cosine with warmup
Training iterations 15k
Context length 2048
Optimizer AdamW
Optimizer hyper-parameters $\beta_{1},\beta_{2},\epsilon$ = 0.9, 0.995, 1e-6
Data MMC4, LAION-2B, LAION-COCO, CC-12M, Objects365
Augmentation CenterCrop
Batch size (per GPU) 2,2,4
Table 11: Hyper-parameters for pre-training.
B.2 Supervised Fine-tuning
VQA and Image Captioning. For this task, we train our model in the form of question answering, i.e., using Based on the image, please answer the question. {image}{question}. The answer is: {answer} as the instruction prompt. We utilize public available datasets for supervised fine-tuning, including LLaVA-Mix-665K [52], COCO Caption [11], VQAv2 [27], ChartQA [60], DocVQA [13], EST-VQA [92], InfoVQA [61], STVQA [92], TextCaps [79], LLaVAR [102], OCR-VQA [62], and DVQA [41]. See Tab. 12 for more training details.
Hyper-parameters Value
Input image resolution $448\times 448$
Learning rate 1e-6 (language model)
1e-5 (others)
Weight decay 0.05
Warmup steps 500
Learning rate schedule cosine
Training iterations 10k
Optimizer AdamW
Optimizer hyper-parameters $\beta_{1},\beta_{2},\epsilon$ = 0.9, 0.999, 1e-8
Batch size 256
Table 12: Hyper-parameters for VQA and image captioning.
Referring Expression Comprehension. Following previous works [10, 6], we train our model in the form of VQA, i.e., the prompt being Question: Provide the bounding box coordinate of the region this sentence describes: {object}, Answer: (x1,y1)(x2,y2). The generated bounding box is considered correct if its intersection over union (IoU) with the GT box is greater than 0.5. Only public available datasets, including datasets from RefCOCO [42], RefCOCO+ [58], and RefCOCOg [58] are utilized to train MM-Interleaved. See Tab. 13 for fine-tuning hyper-parameters.
Hyper-parameters Value
Input image resolution $224\times 224$
Learning rate 2e-5
Weight decay 0.05
Warmup steps 500
Learning rate schedule cosine
Training iterations 10k
Optimizer AdamW
Optimizer hyper-parameters $\beta_{1},\beta_{2},\epsilon$ = 0.9, 0.999, 1e-8
Augmentation RandomHorizontalFlip
Batch size 256
Table 13: Hyper-parameters for referring expression comprehension.
Segmentation-to-Image Translation. Following ControlNet [100], we use BLIP [47] to generate text captions for each image in the ADE20K dataset. We train our model on the training set and evaluate on the validation set. The input sequence is formulated as {segmentation image}{caption}{ground truth image} for each segmentation-image pair. See Tab. 14 for more training details.
Hyper-parameters ADE20K
Input image resolution $224\times 224$
Output image resolution $512\times 512$
Learning rate 1e-5 (image decoder) 1e-4 (others)
Weight decay 0.05
Warmup steps 100
Learning rate schedule cosine
Training iterations 4k
Optimizer AdamW
Optimizer hyper-parameters $\beta_{1},\beta_{2},\epsilon$ = 0.9, 0.98, 1e-5
Augmentation RandomHorizontalFlip
Batch size 512
Table 14: Hyper-parameters for Segmentation-to-Image Translation.
Visual Storytelling. Following previous works [43, 104, 66], MM-Interleaved is finetuned on the following three visual storytelling datasets respectively:
  • VIST [34] is a real-world vision-language dataset, containing 34k and 5k samples for training and evaluation. Each sample is a sequence consisting of 5 text captions and images. During training, we concatenate all the texts and images sequentially and the model is trained to predict all images. During inference, we test the model on generating the last image in the sequence, conditioned on all preceding images and texts following [43, 104]. The evaluation metrics are FID [31] and the CLIP similarity [70] between the generated images and the corresponding real images.
  • PororoSV [49] and FlintstonesSV [56] are two cartoon storytelling datasets, containing 10191/2334/2208 and 20132/2071/2309 samples of the train, validation, and test set, respectively. Each sample is a sequence consisting of 5 text captions and frame images. During training, all the texts and images are concatenated sequentially and the model is trained to predict all images. During inference, the last 4 images are generated auto-regressively given the first image and all preceding captions as condition. FID [31] is used as the evaluation metric.
The finetuning hyper-parameters are listed in Tab. 15.
Hyper-parameters VIST / PororoSV / FlintstonesSV
Input image resolution $224\times 224$
Output image resolution $512\times 512$
Learning rate 1e-4 (image decoder) 1e-5 (others)
Weight decay 0.05
Warmup steps 200
Learning rate schedule cosine
Training iterations 4k
Optimizer AdamW
Optimizer hyper-parameters $\beta_{1},\beta_{2},\epsilon$ = 0.9, 0.98, 1e-8
Augmentation CenterCrop
Batch size 128
Table 15: Hyper-parameters for visual storytelling.
B.3 Evaluation
Dataset Task Split Metric
Caption. COCO [11] Scene description test CIDEr($\uparrow$[87]
Flickr30k [68] Scene description test CIDEr($\uparrow$[87]
NoCaps [2] Scene description test CIDEr($\uparrow$[87]
Image2Paragraph [44] Scene description test CIDEr($\uparrow$[87]
VQA. VQAv2 [27] Scene understanding QA test-dev VQA Acc($\uparrow$[4]
OKVQA [59] External knowledge QA val VQA Acc($\uparrow$[4]
GQA [35] Scene understanding QA test-dev VQA Acc($\uparrow$[4]
VizWiz [29] Scene understanding QA test-dev VQA Acc($\uparrow$[4]
TextVQA [80] Text reading QA val VQA Acc($\uparrow$[4]
VisDial [15] Image dialogue val NDCG($\uparrow$)
REC. RefCOCO [42] Referring experssion comprehension - IoU Acc($\uparrow$)
RefCOCO+ [58] Referring experssion comprehension - IoU Acc($\uparrow$)
RefCOCOg [58] Referring experssion comprehension - IoU Acc($\uparrow$)
Generation. MS-COCO [51] Text-to-image generation val-30K FID($\downarrow$[31]
LN-COCO [69] Text-to-image generation val-30k FID($\downarrow$[31]
ADE20k [106] Segmentation-to-image generation val mIoU($\uparrow$)
VIST [34] Interleaved-context image generation val CLIP-Sim($\uparrow$[70], FID($\downarrow$[31]
PororoSV [49] Interleaved-context multi-image generation test FID($\downarrow$[31]
FlintstonesSV [28] Interleaved-context multi-image generation test FID($\downarrow$[31]
Table 16: Summary of evaluation benchmarks, including image caption, visual question answering, referring experssion comprehension, and image generation.
Benchmarks. Evaluating MM-Interleaved comprehensively requires various benchmarks and datasets, such as image caption, visual question answering, text-to-image generation and so on. All these evaluation tasks and metrics are listed in Tab. 16.
Image Generation. For all image generation tasks, the scale of classifier-free guidance [32] and the total inference step is set as 3.5 and 250 by default.
Text Generation. The prompt templates for each text generation tasks are listed in Tab. 17. The ‘Image Caption (short)’ task includes COCO Caption [11], Flicker30k [68] and NoCaps [2], while the ‘Image Caption (long)’ task includes Image2Paragraph [44].
Task
Prompt Template
Zero-shot
Image Caption (short)
{image} a photo of
Image Caption (long)
{image} Please describe the image in detail. The image depicts
VQA (except VizWiz)
Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz QA
Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
Visual Dialog
[…]
Supervised Fine-tuning
Image Caption (short)
{image} Provide a one-sentence caption for the provided image.
Image Caption (long)
{image} Please describe the image in detail.
VQA (except VizWiz)
Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz VQA
Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
Visual Dialog
[…]
Table 17: Prompt templates for text generation.
Appendix C Qualitative Results
C.1 VL Comprehension
Text Reading QA. As is shown in Fig. 5, MM-Interleaved with MMFS provides more accurate answers when requiring fine-grained details for generating text outputs given VL inputs.
Figure 5: Qualitative results on TextVQA [80]. Each example consists of the user query, the answer given by MM-Interleaved with MMFS, and the answer given by MM-Interleaved without MMFS. The image shapes are normalized for visualization.
Referring Expression Comprehension. The visualization of MM-Interleaved on REC tasks is shown in Fig. 6. Our model with MMFS is also capable of generating more accurate coordinates given the referring expression and the query image.
Figure 6: Referring Expression Comprehension on RefCOCOg [58]. Each example consists of the user query, the box predicted with MMFS, and the box predicted without MMFS. The image shapes are normalized for visualization.
C.2 Image Generation
Segmentation-to-image Translation. Fig. 7 shows the visualization results of MM-Interleaved for segmentation-to-image translation. Given the text prompt and segmentation map, the spatial layout of images generated with MMFS is significantly closer to the original ground-truth image, compared to the baseline results without MMFS.
Figure 7: Segmentation-to-Image Generation on ADE20k [106]. Each row is an example consisting of four images, which are the input segmentation map, the ground truth image, the generated image with MMFS, and the generated image without MMFS respectively. The shape of ground-truth images and segmentation maps is normalized for visualization. When without MMFS, the generated results lack spatial alignment with the input segmentation maps.
Multi-image Generation. In Fig. 8, we compare the multiple images sequentially generated by MM-Interleaved with and without MMFS. The images generated with MMFS achieves better spatial consistency (e.g. the background environment, change of viewpoint, character position relationship etc.) and closer semantic alignment with the interleaved image-text context.
Generating Interleaved Image and Texts. Moreover, the model is further finetuned to generate both images and texts simultaneously for visual storytelling tasks. As is shown in Fig. 9, given the first frame image and caption as context, MM-Interleaved with MMFS generates the following interleaved images and texts coherently, achieving balance between the generation diversity and spatial semantic consistency.
Figure 8: Multi-image Generation on PororoSV [49] and FlintstonesSV [28]. Each example consists of four rows. The first row is the first frame image and all corresponding captions. The second row contains the ground truth images of following frames; The third row are the generated results with MMFS; And the last row are the results generated without MMFS. When without MMFS, the generated multiple images lack content consistency in terms of characters, backgrounds, objects, etc.
Figure 9: Interleaved Image-Text Generation on PororoSV [49] and FlintstonesSV [28]. Each example consists of three columns. The first column is the ground-truth images and captions of all frames. The second column are the generated results with MMFS; And the last column are the results generated without MMFS. Only the caption and image of the first frame is given as condition during generation.
Appendix D Future Work
Future works of MM-Interleaved aims to further scale up both the model and data size, enabling end-to-end training with full-parameters from scratch. In addition, we believe that it is also worth exploring to build a comprehensive evaluation benchmark specifically designed for interleaved image-text modeling.

### v2

@@
-MM-Interleaved: Interleaved Image-Text Generative Modeling via Multi-modal Feature Synchronizer
+
 Abstract
-* Equal contribution.† This work is done when Changyao Tian, Weiyun Wang, and Zhe Chen are interns at Shanghai AI Laboratory.🖂 Corresponding author (daijifeng@tsinghua.edu.cn).
-Developing generative models for interleaved image-text data has both research and practical value.
-It requires models to understand the interleaved sequences and subsequently generate images and text.
+* Equal contribution; 🖂 Corresponding author (daijifeng@tsinghua.edu.cn).† This work is done when Changyao Tian, Weiyun Wang, and Zhe Chen are interns at Shanghai AI Laboratory.
+Developing generative models for interleaved image-text data holds both research and practical value.
+It necessitates models to comprehend interleaved sequences and subsequently generate images and text.
 However, existing attempts are limited by the issue that the fixed number of visual tokens cannot efficiently capture image details, which is particularly problematic in the multi-image scenarios.
 To address this, this paper presents MM-Interleaved, an end-to-end generative model for interleaved image-text data.
-It introduces a multi-scale and multi-image feature synchronizer module, allowing direct access to fine-grained image features in the previous context during the generation process.
+It introduces a multi-scale and multi-image feature synchronizer module (MMFS), enabling intermediately direct access to fine-grained image features from the previous context during the generation process.
 MM-Interleaved is end-to-end pre-trained on both paired and interleaved image-text corpora.
@@
 Code and models are available at https://github.com/OpenGVLab/MM-Interleaved.
+Keywords:
+Interleaved image-text modeling Large language models
 1 Introduction
-Interleaved image-text data, where the text and images are related in content and intertwined in layout, is very common on the internet.
-As an extension of image-text pairs widely used in previous multi-modal models [37, 21, 82, 22, 70], the interleaved format not only covers a broader range of data but also presents more complex information structures.
-Developing multi-modal models that can simultaneously understand and generate such interleaved image-text data has significant research value and practical application potential.
-It broadens the application scope of previous multi-modal models [83] and enables more unified processing of multi-modal data, bringing previously isolated research fields such as text-to-image generation [74, 71] and visual question-answering [48, 53] together.
-Figure 1: Illustration of multi-modal feature synchronizer.
-In the auto-regressive generation of interleaved image-text sequences, besides interacting with low-resolution image tokens with self-attention, tokens in MM-Interleaved could also use the multi-modal feature synchronizer to cross-attend multi-scale high-resolution image features.
-The multi-modal feature synchronizer ensures the cross attention will have exactly the same causal relation between image and text tokens.
+Interleaved image-text data like news and blogs is a sequence of multiple images interspersed with text, which is ubiquitous on the internet.
+As an extension of image-text pairs widely used in previous multi-modal models [37, 21, 83, 20, 71], the interleaved format not only covers a broader range of data but also presents longer and more complex article structures.
+Developing multi-modal models capable of simultaneously comprehending and generating such interleaved image-text data has significant research value and practical application potential.
+It expands the scope of previous multi-modal models [84] and enables more unified processing of multi-modal data, bridging previously disparate research fields such as text-to-image generation [74, 73] and visual question-answering [47, 54].
 With recent progress in multi-modal modeling with Large Language Models (LLMs) [7, 65], exploring how LLMs can be leveraged for interleaved image-text modeling has become increasingly attractive.
@@
 In most multi-modal LLMs, images are encoded as visual tokens from an image tokenizer (also known as image encoder) and fed into LLMs together with text tokens [43, 17, 84, 40].
-To reduce computational demands and the required context length of LLMs, Perceiver Resamplers [3] are commonly used to map each image from up to 1024 tokens to a fixed number of visual tokens (e.g., 32 or 64 tokens) as in Fig.
+However, due to computational and memory constraints, multi-modal LLMs have limited context window sizes (e.g., 2048 or 4096 tokens).
+To reduce computational demands and the required context length of LLMs, Perceiver Resamplers [3] are commonly used to map each image from up to 1024 tokens to a small fixed number of visual tokens (e.g., 32 or 64 tokens) as in Fig.
 2.
-Due to the relatively small token number, critical image details may be overlooked, especially in tasks requiring fine-grained observation.
-Increasing the number of visual tokens [55, 99] can address this issue to some extent.
-While the context length of LLMs is typically limited, allocating a large number (e.g., 2890 in [84]) of visual tokens per image would pose a significant limitation in multi-image scenarios.
-We note that the main problem is that the Perceiver Resampler only takes image features as input and is context-insensitive.
-It compresses each input image to a fixed number of tokens, which makes it harder or even impossible to preserve all information required and fulfill all needs for subsequent generation.
-In contrast, allowing intermediate layers in decoder networks to dynamically extract a small amount of relevant information from the images can compensate for this information loss.
-Such a dynamic extraction mechanism should also efficiently handle scenarios requiring multiple images and high-resolution feature maps.
-Motivated by this, we designed a fine-grained multi-modal feature synchronizer based on the deformable sparse attention [110].
-As shown in Fig.
-1, it enables the multi-modal LLMs to observe multi-scale high-resolution image features from multiple images efficiently.
-This mechanism facilitates the extraction of fine-grained image details with only a small number of input visual tokens, which is particularly applicable to interleaved image-text scenarios with multiple images.
-Based on this module, we propose MM-Interleaved, a novel end-to-end generation model for processing interleaved image-text data, as shown in Fig.
-3.
-Images and text are first mapped into tokens through their respective tokenizers (i.e., image encoder and word embedding) and then fed to the LLM, arranged in their original order.
-A special token <BoI> is introduced to represent “Begin of Image”.
-When the LLM processes image and text sequences, each token in the intermediate layers directly observes multi-image and multi-scale feature maps from the previous context through the proposed feature synchronizer module.
-After being processed by the LLM, the features of the text tokens predict the next text word.
-When the <BoI> token is predicted, an image decoder based on diffusion models is used to predict the next image.
-All previous LLM output features are passed into the image decoder to generate images.
-Each pixel in the intermediate layers of the decoder can also extract details of the previous image through the proposed synchronizer module.
-Our models are pre-trained on a mixture of image-text pairs and interleaved image-text sequences.
-We do not use any in-house data.
-Similar to previous multi-modal LLMs, supervised fine-tuning can further enhance our model capabilities.
-Thanks to end-to-end generative modeling, fine-tuning can be applied to both text and image generation.
-Our model is fine-tuned on several tasks, including visual question-answering, image caption, referring expression grounding, text-to-image generation, segmentation-to-image translation, and visual storytelling.
-Compared with previous image-to-text and text-to-image methods, our model achieves competitive results and shows higher token efficiency.
-Compared with joint image and text generation models, we set the new SotA results for a wide range of tasks.
-Overall, MM-Interleaved can handle interleaved image-text inputs and recognize fine-grained details across multiple images to produce accurate textual descriptions and visually consistent images, demonstrating the potential of high-performance multi-modal models that can understand and generate interleaved image-text content.
-Figure 2: Comparison of different types of image-text generative modeling.
+Due to the relatively small visual token number, critical image details may be overlooked, especially in tasks requiring fine-grained observation.
+Although increasing the number of visual tokens [56, 101] may address this issue to some extent, the token number per image must be restricted to ensure complete input into the LLMs.
+Allocating a large number (e.g., 2890 in [85]) of visual tokens per image would pose a significant limitation especially in multi-image scenarios.
+Figure 1: Illustration of Multi-Modal Feature Synchronizer (MMFS).
+In the auto-regressive generation of interleaved image-text sequences, besides interacting with low-resolution image tokens via self-attention, tokens in MM-Interleaved could also use MMFS to cross-attend multi-scale high-resolution image features.
+MMFS ensures the attention will have exactly the same causal relation between images and text tokens.
+We note that the main problem here is the context insensitivity of Perceiver Resamplers, as it only takes image features as input.
+Each input image is compressed into a fixed number of tokens, making it challenging or even infeasible to preserve all requisite information and accommodate subsequent generation requirements.
+However, such information loss can actually be mitigated through further feature extraction in the intermediate layers of the LLMs.
+Based on the intermediate features of the LLMs, relevant information can be dynamically extracted from the original images to fulfill the perceptual requirements.
+Such a dynamic extraction mechanism can also help to effectively handle scenarios involving multiple images and high-resolution image feature maps.
+Figure 2: Different types of image-text generative modeling.
 (a) and (b) can only generate text.
 (c) and (d) can generate both images and text.
-All types except (d) are limited by the fixed number of visual tokens extracted by context-insensitive Resampler, which can not efficiently capture image details and is problematic in multi-image scenarios.
+All types except (d) are limited by the fixed number of visual tokens extracted by context-insensitive Resampler, which will lose image details and is problematic in multi-image scenarios.
+To facilitate dynamic image feature extraction in the intermediate layers of the LLMs, we design a fine-grained Multi-Modal Feature Synchronizer (MMFS) based on the Deformable Sparse Attention [112], which can efficiently extract relevant information from multi-scale image feature maps and multiple images as shown in Fig.
+1.
+Based on MMFS, we further propose MM-Interleaved, a novel end-to-end generation model for processing interleaved image-text data, as shown in Fig.
+4, which can generate both textual and visual response given interleaved image-text inputs.
+Specifically, input images and text are first mapped into tokens through their respective tokenizers (i.e., image encoder and word embedding) and then fed to the LLM, arranged in their original order.
+A special token <BoI> is introduced to represent “Begin of Image”.
+When the LLM processes the input sequences, each token in the intermediate layers directly observes multi-image and multi-scale feature maps from the previous context through MMFS.
+After being processed by the LLM, the features of the text tokens are then used to predict the next text word.
+When the <BoI> token is predicted, an image decoder based on diffusion models is used to predict the next image.
+All previous LLM output features are passed into the image decoder as the generation conditions.
+With MMFS, the image decoder can also extract details of previous images intermediately.
+MM-Interleaved is pre-trained on a mixture of image-text pairs and interleaved image-text sequences without using any in-house data.
+Similar to previous multi-modal LLMs, supervised fine-tuning can further enhance the model capabilities.
+Due to the end-to-end generative modeling, fine-tuning can be applied to both text and image generation.
+Our model is fine-tuned on several tasks, including visual question-answering, image captioning, referring expression grounding, text-to-image generation, segmentation-to-image translation, and visual storytelling.
+As is illustrated in Fig.
+3, our model achieves competitive results and shows higher token efficiency compared with previous image-to-text and text-to-image methods.
+Compared with joint image and text generation models, we set the new SOTA results for a wide range of tasks.
+The key contributions are:
+  • We propose Multi-Modal Feature Synchronizer (MMFS) to reduce the number of visual tokens required by multi-modal LLMs, which can efficiently extract fine-grained visual details from multi-scale feature maps of multiple images according to the intermediate context features of the multi-modal LLMs;
+  • We propose MM-Interleaved based on MMFS for generative modeling of interleaved image-text data, which can preserve fine-grained image information with only a small number of visual tokens per image, and can be optimized end-to-end for both text and image generation;
+  • Our method can generate both accurate text descriptions and visually consistent images given interleaved image-text inputs, achieving SOTA results on various multi-modal comprehension and generation benchmarks without using any in-house data.
+Figure 3: Comparison on multi-modal benchmarks.
+11 representative tasks are listed.
+See Sec.
+4 for other benchmarks.
+MM-Interleaved (red) consistently outperforms previous methods supporting interleaved generation (green).
+It also surpasses or is on par with specialists in either text generation (brown) or image generation (blue).
 2 Related Work
 Modeling of Paired Image-Text Data plays an important role in the progress of multi-modal research in recent years.
-[37, 24, 76] has released a series of public large-scale image-text pairs datasets.
-Based on these data, models trained by image-text contrastive learning [70, 39, 82, 50] have been proven able to recognize and understand open-world semantics.
-Subsequent work [93, 111, 46, 97] further incorporates the text generation tasks such as image captioning.
-Some other works [74, 71] were proposed to generate images based on text prompts.
-The latest progress of LLMs [64, 65] has created a new era, leading to the emergence of many LLM-based multi-modal models trained using image-text pairs [48, 53, 90, 91].
-For example, [53] connect pre-trained LLMs and vision foundation models with linear projections to build powerful multi-modal text-generation models.
-Figure 3: Architecture of the proposed MM-Interleaved Model.
-The red lines represent how the multi-scale image features are generated and utilized.
-MM-Interleaved incorporates image encoder to both extract high-resolution multi-scale image features and map each image into a fixed number of low-resolution visual tokens.
-These visual tokens are fed into the multi-modal LLM along with text tokens.
-LLM uses the feature synchronization module to extract high resolution image details, and auto-regressively generates text tokens.
-After that, a DM-based image decoder generates next image conditioned on the previous context features from LLM as well as the multi-scale image features from the feature synchronization module.
+A series of public large-scale image-text pairs datasets have been released [37, 23, 77].
+Based on these data, models trained by image-text contrastive learning [71, 39, 83, 50] are able to recognize and understand open-world semantics.
+Subsequent works [94, 113, 49, 99] further incorporated text generation tasks such as image captioning, while other works [74, 73] were proposed to generate images based on text prompts.
+The latest progress of LLMs [65, 66] has launched a new era, leading to the emergence of many LLM-based multi-modal models trained using image-text pairs [47, 54, 92, 91].
+For example, LLaVA [54] connects LLMs with pretrained visual encoders using linear projections to build powerful multi-modal foundation models.
 Modeling of Interleaved Image-Text Data has received increasing attention recently.
-[33, 3] were the first to focus on understanding such data relying on non-public datasets.
-To promote the development of this field, [109, 45] release public and large-scale interleaved image-text datasets.
-More works [103, 101] have begun to focus on understanding the interleaved data.
-Nevertheless, their generative capabilities remain limited to text.
-[43, 99, 83, 18] have initially attempted to generate images and text in the interleaved context concurrently.
-[43, 83] involve a two-stage generation process, where text and image are only generated in the first and second stages, respectively.
-[99] uses VQ-VAE [86] to convert images into discrete tokens, facilitating token-level auto-regressive modeling as traditional language modeling.
+Early works such as Kosmos [33] and Flamingo [3] focused on understanding such data with non-public datasets.
+To promote the development of this field, public and large-scale interleaved image-text datasets were later released in [111, 45].
+More recent works [105, 103] have concentrated on understanding the interleaved data.
+Nevertheless, their generative capabilities remain confined to text.
+Initial attempts at generating images and text given interleaved contexts were first undertaken in [43, 101, 84, 17].
+A two-stage generation process was introduced by [43, 84], wherein text and images are generated in the first and second stages, respectively.
+CM3Leon [101] employs VQ-VAE [87] to convert images into discrete tokens, facilitating token-level auto-regressive modeling as language modeling.
 However, it primarily focuses on image generation capabilities and exhibits notable weaknesses in image understanding.
-DreamLLM [18], a concurrent work to ours, also focuses on single-stage end-to-end modeling using raw image pixels as input.
-Despite these efforts, they only feed image information at the input of LLMs, which are limited by the problem that fixed number of visual tokens cannot efficiently describe image details.
+DreamLLM [17] focuses on single-stage end-to-end modeling using raw image pixels as inputs.
+Emu2 [82], SEED-LLaMA [25] and VL-GPT [110] adopt additional training stage for the image tokenizer-detokenizer.
+Despite these efforts, they only feed image information at the LLM inputs, which are limited by the problem that fixed number of visual tokens cannot efficiently preserve image details.
 Integrating Image Details into LLMs is important for multi-modal LLMs.
 Most works use Perceiver Resamplers [3, 47, 109] to extract image information via cross-attention, mapping each image into a fixed number of visual tokens.
-For example, [3] equips Resamplers in the intermediate layers of LLMs, injecting extracted image features into LLMs through gated residual connections.
-[48, 107] introduces Resamplers at the bottom of LLM to insert the extracted visual tokens into the input text sequences.
-While these methods achieve good performance, image details may be overlooked due to the small number (e.g., 32 or 64) of visual tokens.
-To preserve more image details, [6, 55, 53, 10] increase the number of input visual tokens per image to hundreds.
-[84] further increases this number to 2,890.
-Although it mitigates information loss, the computational and memory demands of LLMs are significantly increased.
+For example, Flamingo [3] adopts Resamplers in the intermediate layers of LLMs, injecting extracted image features into LLMs through gated residual connections.
+BLIP-2[47] and Mini-GPT4 [109] introduce Resamplers at the bottom of LLM to insert the extracted visual tokens into the input text sequences.
+While these methods achieve good performance, image details remain to be overlooked due to the small number (e.g., 32 or 64) of visual tokens.
+To preserve more details, recent works [6, 56, 54, 10] increase the number of input visual tokens per image to hundreds.
+In SPHINX [85], the token number is further increased to 2,890.
+Although it helps mitigate information loss, the computational and memory demands of LLMs are also significantly increased.
 Increasing the number of visual tokens per image is particularly problematic in multi-image scenarios, where multiple images naturally require more visual tokens, making it hard to use for interleaved image-text data.
@@
 3.1 Task Formulation
-To build an end-to-end generative model for interleaved image-text data.
-We first consider an interleaved image-text sequence $X=\{x_{1},x_{2},x_{3},\dots\}$, where each element $x_{n}$ is either a text token (denoted as $x_{n}^{L}$) or a whole image (denoted as $x_{n}^{V}$).
+To build an end-to-end generative model for interleaved image-text data, we first consider an interleaved image-text sequence $X=\{x_{1},x_{2},x_{3},\dots\}$, where each element $x_{n}$ is either a text token (denoted as $x_{n}^{L}$) or a whole image (denoted as $x_{n}^{V}$).
 Text and images are arranged in the order in which they appear in the original content.
-In multi-modal LLMs, a common practice is first extracting embedding for each text token and each image before being fed into LLMs, i.e., $e_{n}^{L}=\mathcal{E}_{L}(x_{n}^{L})$ and $e_{n}^{V}=\mathcal{E}_{V}(x_{n}^{V})$, where $\mathcal{E}_{L}$ denotes word embedding following standard practices in NLP.
+In multi-modal LLMs, a common practice is to first extract embeddings for each text token and each image and then feed them into LLMs, i.e., $e_{n}^{L}=\mathcal{E}_{L}(x_{n}^{L})$ and $e_{n}^{V}=\mathcal{E}_{V}(x_{n}^{V})$, where $\mathcal{E}_{L}$ denotes word embedding following standard practices in NLP.
 $\mathcal{E}_{V}$ is typically an image encoder (e.g., ViTs [18]) followed by a Perceiver Resampler [3] to map each image to a fixed number of tokens.
-Then, the generative modeling is trained to maximize the log-likelihood of the multi-modal sequence, given by:
-$\log p(X)$ $=\sum_{n}\log p(x_{n}|x_{1},x_{2},\dots,x_{n-1})$ (1)
-$=\sum_{n}\log p(x_{n}|e_{1},e_{2},\dots,e_{n-1})$
-$=\sum_{n\in\mathcal{I}_{L}}\log p(x_{n}^{L}|e_{<n})+\sum_{n\in\mathcal{I}_{V}}\log p(x_{n}^{V}|e_{<n}),$
-where $\mathcal{I}_{L}$ and $\mathcal{I}_{V}$ represent the index sets for text tokens and images within the sequence, respectively.
+Then, the generative modeling is trained to maximize the log-likelihood:
+$\log p(X)=\sum_{n}\log p(x_{n}|e_{<n})=\sum_{n\in\mathcal{I}_{L}}\log p(x_{n}^{L}|e_{<n})+\sum_{n\in\mathcal{I}_{V}}\log p(x_{n}^{V}|e_{<n}),$ (1)
+where $\mathcal{I}_{L}$ and $\mathcal{I}_{V}$ represent the index sets for text tokens and images, respectively.
 $<n$ in the subscript represents the abbreviation of $\{1,2,\dots,n-1\}$.
@@
 (1).
+Figure 4: Architecture of MM-Interleaved.
+The red lines represent how the multi-scale image features are generated and utilized.
+MM-Interleaved incorporates image encoder to both extract high-resolution multi-scale image features (red lines) and map each image into a fixed number of low resolution visual tokens.
+These visual tokens are fed into the multi-modal LLM along with text tokens.
+LLM then uses our proposed feature synchronizer (MMFS) to extract additional high-resolution image details, and auto-regressively generate text tokens.
+After that, a diffusion-based image decoder generates next image conditioned on the previous context features from LLM, where MMFS is also utilized to capture accurate visual conditions.
 Text Generation with Multi-modal Condition.
-The log-likelihood $\log p(x_{n}^{L}|e_{<n})$ is similar to traditional causal language modeling, except that the condition also includes previous images.
+$\log p(x_{n}^{L}|e_{<n})$ is similar to traditional causal language modeling, except that the condition also includes previous images.
 Recent works [47, 54, 92, 91] have demonstrated the effectiveness of using LLMs for processing additional visual inputs.
@@
 The loss function for text generation is
-$L_{\text{NTP}}(x_{n}^{L}|e_{<n})=-\bar{x}_{n}^{L}\cdot\log\mathrm{softmax}\big{(}W\cdot\mathcal{D}_{\text{LM}}(e_{<n})\big{)},$ (2)
-where $W$ is the linear classification weight, $\mathcal{D}_{\text{LM}}$ denotes the LLM network (e.g., LLaMA [85]), $\bar{x}_{n}^{L}$ is the one-hot vector indicating the ground-truth text token.
+$L_{\text{NTP}}(x_{n}^{L}|e_{<n})=-\bar{x}_{n}^{L}\cdot\log\mathrm{softmax}\big{(}W\cdot\mathcal{D}_{\text{LLM}}(e_{<n})\big{)},$ (2)
+where $W$ is the linear classification weight, $\mathcal{D}_{\text{LLM}}$ denotes the LLM network (e.g., LLaMA [86]), $\bar{x}_{n}^{L}$ is the one-hot vector indicating the ground-truth text.
 Image Generation with Multi-modal Condition.
-Maximizing the log-likelihood $\log p(x_{n}^{V}|e_{<n})$ aligns with the denoising-diffusion process [54], which recently achieves widespread success in image generation.
-When conditional inputs $x_{<n}$ are all text, this is identical to text-to-image diffusion models, where maximizing the log-likelihood is derived as minimizing the diffusion modeling loss as
-$L_{\text{NIP}}(x_{n}^{V}|e_{<n})=\mathbb{E}_{\epsilon,t}~{}||\epsilon-\mathcal{D}_{\text{DM}}\big{(}x_{n,t}^{V},t,\mathcal{D}_{\text{LM}}(e_{<n})\big{)}||^{2},$ (3)
-where $\mathcal{D}_{LM}$ is the language model encoding text inputs, $\mathcal{D}_{DM}$ is the diffusion model for image denoisiong.
+Maximizing $\log p(x_{n}^{V}|e_{<n})$ aligns with the denoising-diffusion process [55], which recently achieves widespread success in image generation.
+Maximizing the log-likelihood is derived as minimizing the diffusion modeling loss as
+$L_{\text{NIP}}(x_{n}^{V}|e_{<n})=\mathbb{E}_{\epsilon,t}~{}||\epsilon-\mathcal{D}_{\text{DM}}\big{(}x_{n,t}^{V},t,\mathcal{D}_{\text{LLM}}(e_{<n})\big{)}||^{2},$ (3)
+where $\mathcal{D}_{DM}$ is the diffusion model for image denoisiong.
 $x_{n,t}^{V}$ is the noisy version of the original image at the denoising step $t$, and the denoising network $\mathcal{D}_{DM}$ is trained to predict the noise $\epsilon$.
 We found that such modeling is also applicable when conditional input is interleaved image-text data.
-Note that our modeling allows for the flexible combination of different language models, image encoders, and image decoders to fully leverage a variety of pre-trained models.
-It does not need to insert additional special tokens specifically for image generation into LLMs.
-The whole framework can be optimized end-to-end.
+Such modeling allows for the flexible combination of different language models, image encoders, and image decoders to fully leverage a variety of pre-trained models.
+The entire framework can be optimized end-to-end.
 3.2 Architecture
@@
 It integrates a Visual Foundation Model (VFM), a Large Language Model (LLM), and a Diffusion Model (DM).
-This integration aims to excel in both understanding and generation tasks of text and images by leveraging the strengths of each model type.
+Such integration aims to excel in both comprehension and generation tasks of text and images by leveraging the strengths of each model type.
 As illustrated in Fig.
-3, our architecture comprises three key components:
+4, our architecture comprises three key components:
 (1) VFM-based Image Tokenizer $\mathcal{E}_{V}$ that maps each image $x^{V}\in\mathbb{R}^{H\times W\times 3}$ (e.g., $H=W=224$) into a fixed number of visual tokens $e^{V}\in\mathbb{R}^{N\times C}$ ($N=32$ by default).
 $C$ is the channel dimension.
-It consists of a pre-trained vision foundation model (e.g., CLIP-ViT [70]) for feature extraction and a Perceiver Resampler [3] to reduce the number of visual tokens.
-We also use a ViT-Adapter [12] to extract multi-scale image features $F^{V}\in\mathbb{R}^{(\sum_{i=1}^{L}H_{i}\times W_{i})\times C}$ for fine-grained feature fusion in subsequent networks, where $L=4$ and $H_{i}=\frac{H}{2^{i+1}},W_{i}=\frac{W}{2^{i+1}}$ by default.
-(2) LLM-based Multi-modal Model $\mathcal{D}_{\text{LM}}$ that extracts context features from the interleaved image-text sequences.
+It consists of a pre-trained vision encoder (e.g., CLIP-ViT [71]) for feature extraction and a Perceiver Resampler [3] to reduce the number of visual tokens.
+A ViT-Adapter [12] is also used to extract multi-scale image features $F^{V}\in\mathbb{R}^{(\sum_{i=1}^{L}H_{i}\times W_{i})\times C}$ for fine-grained feature fusion in subsequent networks, where $L=3$ and $H_{i}=\frac{H}{2^{i+2}},W_{i}=\frac{W}{2^{i+2}}$ by default.
+(2) LLM-based Multi-modal Model $\mathcal{D}_{\text{LLM}}$ that extracts context features from the interleaved image-text sequences.
 A pre-trained LLM (e.g., Vicuna [107]) is utilized.
-Its input sequence $E\in\mathbb{R}^{K\times C}$ is a concanation of embeddings $(e_{1},e_{2},\dots)$, where $e_{n}$ is either a word embedding $e_{n}^{L}\in\mathbb{R}^{1\times C}$ or an image embedding $e_{n}^{V}\in\mathbb{R}^{N\times C}$.
+Its input sequence $E\in\mathbb{R}^{K\times C}$ is a concatenation of embeddings $(e_{1},e_{2},\dots)$, where $e_{n}$ is either a word embedding $e_{n}^{L}\in\mathbb{R}^{1\times C}$ or an image embedding $e_{n}^{V}\in\mathbb{R}^{N\times C}$.
 $K$ is the total number of input tokens.
-We also introduce a fine-grained feature synchronizer module for letting the intermediate layers in $\mathcal{D}_{\text{LM}}$ directly access and extract multi-scale image features on demand.
+We also introduce a feature synchronizer (i.e., MMFS) for letting the intermediate layers in $\mathcal{D}_{\text{LLM}}$ directly access and extract multi-scale image features on demand.
 (3) DM-based Image Decoder $\mathcal{D}_{\text{DM}}$ that generates the image based on previous image-text sequences.
@@
 Then we introduce the details of the proposed architecture.
-Figure 4: Architecture of the MMFS module.
-The query feature is passed through a linear layer and added to the image index embedding.
-Two linear layers are used to predict the sampling offsets and unnormalized attention weights of each image respectively.
-The sampling offsets are added with the query’s reference point to form the corresponding sampling locations, which are shared and rescaled across all feature maps of the same image.
-The output feature is the weighted sum of the sampled spatial features from these points.
-Multi-image and Multi-scale Feature Synchronizer (MMFS).
-The MMFS module is designed to dynamically capture image details from multiple images in an interleaved context.
-It leverages the Deformable Attention [110] mechanism to achieve efficient and sparse image attention.
+Multi-Modal Feature Synchronizer (MMFS).
+MMFS aims to enable dynamic and efficient image detail extraction in decoder intermediate layers, compensating for information loss due to limited input visual tokens.
+It leverages the Deformable Attention [112] to achieve efficient and sparse image attention.
+MMFS can be applied to both image and text decoding, avoiding information bottlenecks caused by the Resamplers in multi-modal LLMs.
+It is especially efficient for processing multiple high-resolution images in context.
+Figure 5: Architecture of MMFS module.
+The query feature is passed through a linear projection and added with the image index embedding.
+Two linear projections are used to predict the sampling offsets and unnormalized attention weights of each image, respectively.
+The sampling offsets are added with the query’s reference point to form the corresponding sampling locations, which are shared across all feature maps of the same image.
+The output is the weighted sum of the sampled spatial features.
 As is shown in Fig.
-4, given a query token that requires detailed image features, the MMFS module only attends to a small set of sampling points around a reference point on the reference images.
-Let $f_{q}\in\mathbb{R}^{C}$ represent the features of the query token.
-$\hat{p}_{q}\in[0,1]^{2}$ denotes the relative image coordinate of its reference point.
-By default, $\hat{p}_{q}=(0.5,0.5)$ corresponds to the center of the images.
-$\{F^{V}_{m}\}_{m=1}^{M}$ are the multi-image multi-scale feature maps, where $F^{V}$ is the multi-scale feature extracted by the image tokenizer, $M$ is the number of reference images.
+5, given a query token that requires detailed image features, the MMFS module only attends to a small set of sampling points around a reference point on the reference images.
+Let $f_{q}\in\mathbb{R}^{C}$ represents the features of the query token.
+$\hat{p}_{q}\in[0,1]^{2}$ denotes the relative image coordinate of its reference point, which is designed for compressing the search space for sampling points from global positions to local relative offsets.
+By default, without a spatial prior, $\hat{p}_{q}=(0.5,0.5)$ is set as the center of the images.
+$\{F^{V}_{m}\}_{m=1}^{M}$ are the multi-image multi-scale feature maps extracted by the image tokenizer, $M$ is the number of reference images.
 The output feature $f_{o}\in\mathbb{R}^{C}$ is given by
@@
 $p_{q}^{(m)}$ $=W_{p}\cdot q^{(m)}+\hat{p}_{q},\ \ A_{q}^{(m)}=W_{A}\cdot q^{(m)},$
-$p_{q}$ $=\mathrm{Concat}(p_{q}^{(1)},\cdots,p_{q}^{(M)}).$
-$A_{q}$ $=\mathrm{softmax}\left(\mathrm{Concat}(A_{q}^{(1)},\cdots,A_{q}^{(M)})\right).$
+$p_{q}$ $=\mathrm{Concat}(p_{q}^{(1)},\cdots,p_{q}^{(M)}),$
+$A_{q}$ $=\mathrm{softmax}\left(\mathrm{Concat}(A_{q}^{(1)},\cdots,A_{q}^{(M)})\right),$
 $f_{o}$ $=\mathrm{DeformAttn}(\{F^{V}\}_{m=1}^{M},A_{q},p_{q}),$
@@
 Here, $L$ and $K$ denote the number of multi-scale feature levels and sampling points per feature map, respectively.
-MMFS can be applied to both image and text decoding, avoiding information bottlenecks caused by the widely used Resamplers in multi-modal LLMs.
-It is especially efficient for processing multiple high-resolution images in context.
 Multi-modal LLM with MMFS.
@@
 Image token embeddings are inserted at the corresponding positions in the original sequence.
-Another special token <BoI> is added in front of each image to represent “Begin of Image”.
+Special token <BoI> is added in front of each image to represent “Begin of Image”.
 MMFS modules are inserted between the self-attention layer and the feed-forward layer of the LLM every fixed number of blocks.
@@
 $\hat{p}_{q}=(0.5,0.5)$.
-The output of MMFS is multiplied by tanh($\alpha$) before being added back to $f_{q}$, where $\alpha$ is a zero-initialized learnable scalar.
+The output of MMFS is multiplied by tanh($\alpha$) before added back to $f_{q}$, where $\alpha$ is a zero-initialized learnable scalar.
 Image Decoder with MMFS.
-The output features from the multi-modal LLM are further processed by another Perceiver Resampler before being fed into the image decoder.
-For each image to be generated, the Resampler maps previous context features to a fixed number of tokens (e.g., 77 tokens) to match the pre-trained diffusion model.
-MMFS modules are inserted after each upsampling block in the U-Net structure of the diffusion model.
+The output features from the multi-modal LLM are further processed by another Resampler before fed into the image decoder.
+For each image to be generated, the Resampler maps previous context features to a fixed number of tokens (e.g., 77 tokens) to match the condition token length of the pre-trained diffusion model.
+MMFS modules are inserted after each downsampling block in the U-Net of the diffusion model.
 Query token $f_{q}$ iterates over each pixel in the feature map.
 $\hat{p}_{q}$ is set as the spatial coordinate of the query pixel.
-The output of MMFS is further processed by a zero-initialized convolution layer before being added back to $f_{q}$.
-Training Target and Inference Pipeline. The training objective is defined as the sum of Next-Text-Token Prediction loss in Eq. (2) and Next-Image Prediction loss in Eq. (3),
-$\mathcal{L}=\mathcal{L}_{NTP}+\lambda~{}\mathcal{L}_{NIP}$ (5)
-where $\lambda$ is a hyperparameter used to determine the relative loss weight between the image and text decoding branches. The whole framework can be optimized end-to-end.
+The output of MMFS is further processed by a zero-initialized convolution before added back to $f_{q}$.
+Training Target and Inference Pipeline.
+The training objective is defined as the sum of Next-Text-Token Prediction loss in Eq.
+(2) and Next-Image Prediction loss in Eq.
+(3) as $\mathcal{L}=\mathcal{L}_{NTP}+\lambda~{}\mathcal{L}_{NIP}$, where $\lambda$ is a hyperparameter used to determine the relative loss weight between the image and text decoding branches.
+The whole framework can be optimized end-to-end.
 During inference, the images and texts are generated in an auto-regressive manner.
@@
 4.1 Implementation Details
-[…]
+Network.
+We adopt CLIP-ViT-L/14 [71], Vicuna-13B [107] and Stable Diffusion v2.1 [74] as the image encoder, large language model, and image decoder, respectively.
+For the multi-modal LLM, a Perceiver Resampler with 12 blocks is used to reduce the number of visual tokens per image to 64.
+MMFS is inserted every 4 blocks in the LLM.
+For the image decoder, a Perceiver Resampler with only 1 block is used to reduce the number of previous conditional tokens to 77.
+MMFS is inserted after each downsampling block in the image decoder.
+Pre-training.
+Our model is pre-trained on a mixture of image-text pairs and interleaved image-text sequences, including MMC4 [111], LAION-2B [77], LAION-COCO [78], CC-12M [8] and Objects365 [79].
+For CC-12M [8] and Objects365 [79], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [47] to caption the images.
+The sampling probability of MMC4 is twice that of other image-text pair datasets.
+No in-house data is used.
+The images are inserted before or after the corresponding text sentence with equal probability.
+To optimize training efficiency and data utility, multiple image-text pairs or interleaved image-text sequences are concatenated into extended sequences with the maximum context length (i.e., 2048 tokens).
+The model is pre-trained for 15000 steps with 11 billion tokens visited.
+The image encoder and LLM are frozen.
+The learning rate is set to be $10^{-5}$ for the image decoder and $10^{-4}$ for the rest trainable parameters.
+The input and output image resolutions are $224\times 224$ and $512\times 512$, respectively.
+Model LLM H I A COCO Flickr NoCaps I2Para. VQAv2 OKVQA GQA VizWiz TextVQA VisDial
 Models for Text-Generation Only
@@
 LLaVA-1.5 [53] Vicuna-13B - I A 80.0 63.3 53.6 61.3
-VL-GPT-I [108] LLaMA-7B - I A 133.7 67.2 50.3 51.5 38.9 51.8
 Qwen-VL [6] Qwen-7B H I A 85.8 121.4 78.8 59.3 35.2 63.8
@@
 VL-GPT [110] LLaMA-7B - - - 116.4 51.7 35.8 34.6 34.7 49.9
-MM-Interleaved Vicuna-13B - - - 129.0 85.8 106.4 23.5 57.0 40.0 40.8 37.2 48.7
+VL-GPT-I [110] LLaMA-7B - I A 133.7 67.2 50.3 51.5 38.9 51.8
+SEED-LLaMA [25] LLaMA2-Chat-13B - I A 125.0 48.1 27.1 23.3
+SEED-LLaMA-I [25] LLaMA2-Chat-13B - I A 126.9 63.4 43.2 49.4
+MM-Interleaved Vicuna-13B - - - 129.0 85.8 106.4 23.5 57.0 40.0 46.7 40.8 37.2 48.7
 MM-Interleaved-SFT Vicuna-13B - I A 140.5 93.0 123.2 30.3 80.2 51.7 60.5 54.9 61.0 53.7
-[…]
-Network.
-We adopt CLIP-ViT-L/14 [70], Vicuna-13B [105] and Stable Diffusion v2.1 [74] as the image encoder, large language model, and image decoder, respectively.
-For the multi-modal LLM, a Perceiver Resampler with 12 blocks is used to reduce the number of visual tokens per image to 64.
-The MMFS module is inserted every 4 blocks in the LLM.
-For the image decoder, a Perceiver Resampler with only 1 block is used to reduce the number of conditional tokens per image to 77.
-The MMFS module is inserted after each up-sampling block in the image decoder.
-Pre-training.
-Our models are pre-trained on a mixture of interleaved image-text sequences and image-text pairs, including MMC4 [109], LAION-2B [76], LAION-COCO [77], CC-12M [8] and Objects365 [78].
-For CC-12M [8] and Objects365 [78], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [48] to caption the images.
-The sampling probability of MMC4 is twice that of other image-text pair datasets.
-No in-house data is used.
-The images are inserted before or after the corresponding text sentence with equal probability.
-To optimize training efficiency and data utility, multiple image-text pairs or interleaved image-text sequences are concatenated into extended sequences with the maximum context length (i.e., 2,048 tokens).
-The model is pre-trained for 15,000 steps with 11 billion tokens visited.
-The image encoder and LLM are frozen.
-The learning rate is set to be $10^{-5}$ for the image decoder and $10^{-4}$ for the rest trainable parameters.
-The input and output image resolutions are $224\times 224$ and $512\times 512$, respectively.
-Supervised Fine-tuning.
-After pre-training, like multi-modal LLMs, our model capabilities can be further enhanced using supervised fine-tuning.
-The model is fine-tuned on four types of downstream tasks, which are 1) visual question-answering and image caption, 2) referring expression comprehension, 3) segmentation-to-image translation, and 4) visual storytelling.
-More details of datasets and hyper-parameters used for pre-training and fine-tuning could be found in the Appendix.
-Model RefCOCO [42] RefCOCO+ [58] RefCOCOg [58]
-Val Test-A Test-B Val Test-A Test-B Val Test
-OFA-L [88] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.50
-VisionLLM-H [90] - 86.70 - - - - - -
-Shikra [10] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
-MiniGPT-V2 [9] 88.69 91.65 85.33 79.97 85.12 74.45 84.44 84.66
-Ferret [96] 89.48 92.41 84.36 82.81 88.14 75.17 85.83 86.34
- Qwen-VL [6] 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
-MM-Interleaved (ours) 89.92 92.59 86.54 82.99 88.57 77.07 85.21 84.92
-Table 2: Supervised fine-tuning results on referring expression comprehension task.
-“*” denotes using an additional self-constructed grounding dataset and trained with an image resolution larger than 224.
+Table 1: Multi-modal comprehension evaluation.
+“H” denotes using in-house data, “I” means the training images of some benchmarks are included in the training, “A” means the training annotations of some benchmarks are visible in training.
+Benchmarks include COCO [11]; Flickr: Flickr30k [69]; NoCaps [2]; I2Para.: Image2Paragraph [44]; VQAv2: VQAv2 [27]; OKVQA [60]; GQA [35]; VizWiz [29]; TextVQA [81]; VisDial [15].
+The results in underline and bold are the best performance without “HIA” data and without “H” data, respectively.
+The evaluation metrics for each benchmark are listed in Appendix.
+Model LLM $L_{1}$ (Part-1) Part-2 $L_{2}$ Part-3 $L_{3}$
+Emu [84] LLaMA-13B 42.5 41.1 42.4 41.4 42.3
+Next-GPT [95] Vicuna-7B 30.7 35.6 31.1 33.9 31.4
+SEED-LLaMA [25] LLaMA2-Chat-13B 43.9 43.4 43.8 52.3 44.8
+MM-Interleaved Vicuna-13B 43.9 46.1 44.1 52.1 45.0
+Table 2: Zero-shot results for interleaved image-text comprehension and generation on SEED-Bench-2 [46].
+The average task accuracy across corresponding evaluation dimensions is reported.
+$L_{1}$(part-1) evaluates the image and text comprehension, $L_{2}$(part-1&2) evaluates interleaved image-text comprehension, and $L_{3}$(part-1&2&3) evaluates image and text generation.
 Model MS-COCO LN-COCO
@@
 DALL-E [73] $\sim$28 -
-CogView [16] 27.10 -
 CogView2 [16] 24.00 -
@@
 Emu-13B [84] 11.66 -
+Next-GPT [95] 11.28 -
 CM3Leon-7B [101] 10.82 -
@@
 DreamLLM-7B [17] 8.46 20.53
-MM-Interleaved-13B (ours) 7.90 23.88
-Table 3: Zero-shot text-to-image generation results on MS-COCO [51] and LN-COCO [69]. FID [31] is reported.
+MM-Interleaved 7.90 23.88
+Table 3: Zero-shot text-to-image generation results. FID [31] is reported.
+Supervised Fine-tuning.
+After pre-training, like other multi-modal LLMs, our model capabilities can be further enhanced using supervised fine-tuning.
+The model is end-to-end fine-tuned on four types of downstream tasks: 1) visual question-answering and image caption, 2) visual storytelling, 3) segmentation-to-image translation, and 4) referring expression comprehension.
+More implementation details could be found in Appendix.
 4.2 Evaluation
-We evaluate the zero-shot visual-language capability of the pre-trained MM-Interleaved as well as its performance on various downstream tasks after fine-tuning and report the results in Tab. 1.
+We evaluate the zero-shot capability of the pre-trained MM-Interleaved as well as its performance on various downstream tasks after supervised fine-tuning.
 Zero-shot Results after Pre-training.
-Our method demonstrates strong multi-modal zero-shot comprehension on various benchmarks, including image captioning on COCO [11], Flicker30K [68], NoCaps [2] and Image2Paragraph [44]; general visual question answering on VQAv2 [27], OKVQA [59], VizWiz [29] and TextVQA [80].
-We also evaluate on the VisDial [15] visual dialogue task.
-In the fully decontaminated setting where the image and text in downstream tasks are unseen during pre-training, our model significantly outperforms MetaLM [30], OF-9B [5] and IDEFICS-80B [45] on all captioning, VQA, and multi-modal tasks.
+Tab.
+1 demonstrates our strong performance on multi-modal zero-shot comprehension across various benchmarks, including image captioning (COCO [11], Flicker30K [69], NoCaps [2], I2Para [44]), visual question answering (VQAv2 [27], OKVQA [60], VizWiz [29], TextVQA [81]) and visual dialogue (VisDial [15]).
+In the fully decontaminated setting where the image and text in downstream tasks are unseen during pre-training, our model significantly outperforms other methods on all tasks.
 This demonstrates the effectiveness of the proposed MM-Interleaved approach.
 Furthermore, our model even exceeds most trained on vast in-house data like Flamingo-9B [3] and Emu [84], showing the importance of proper architecture for image-text interaction.
-Additionally, we evaluate text-conditional image generation on MS-COCO [51] and LN-COCO [69].
+Tab.
+3 shows the results on text-to-image generation for MS-COCO [52] and LN-COCO [70].
 On MS-COCO, we sample 8 images per text condition and use CLIP [71] to rerank based on text-image similarity.
 CLIP reranking is not used for LN-COCO.
-FID [31] is used to evaluate both datasets.
-As show in Tab.
-3, our model shows competitive text-to-image generation compared to existing image and text generation models.
-Fine-tuning Results of VQA and Captioning.
-Our fine-tuned model achieves state-of-the-art performance on image captioning benchmarks.
-On visual question answering tasks, it matches the previous best LLaVA-1.5 model [52].
-Compared to LLaVA-1.5, our model has two key advantages 1) our method is capable of generating both images and text, while LLaVA-1.5 can only generate text; 2) LLaVA-1.5 uses 576 visual tokens as LM inputs, whereas we only require 64.
+Our model achieves competitive performance compared to existing image and text generation models.
+Note that some other works (e.g., Emu, Muse, Imagen, and Parti) use in-house data, while ours did not.
+Tab.
+2 shows the zero-shot performance for interleaved image-text comprehension and generation on SEED-Bench-2 [46].
+Our pre-trained model achieves SOTA results on both comprehension and generation tasks.
+Visualization results for interleaved image-text generation are showed in Appendix.
+Model RefCOCO [42] RefCOCO+ [59] RefCOCOg [59]
+Val Test-A Test-B Val Test-A Test-B Val Test
+OFA-L [89] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.50
+VisionLLM-H [92] - 86.70 - - - - - -
+Shikra [10] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
+MiniGPT-V2 [9] 88.69 91.65 85.33 79.97 85.12 74.45 84.44 84.66
+Ferret [98] 89.48 92.41 84.36 82.81 88.14 75.17 85.83 86.34
+Qwen-VL [6] 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
+MM-Interleaved 89.92 92.59 86.54 82.99 88.57 77.07 85.21 84.92
+Table 4: Supervised fine-tuning results on referring expression comprehension task.
+* denotes using an additional self-constructed grounding dataset and trained with an image resolution larger than 224.
+Fine-tuning Results.
+Tab.
+1 also shows our fine-tuned results (MM-Interleaved-SFT) on multi-modal comprehension.
+Our fine-tuned model achieves SOTA performance without using any in-house data.
+On visual question answering tasks, it matches the previous best LLaVA-1.5 model.
+Compared to LLaVA-1.5, our model has two key advantages: 1) our method is capable of generating both images and text, while LLaVA-1.5 can only generate text; 2) LLaVA-1.5 uses 576 visual tokens as LLM inputs, whereas we only require 64 tokens.
 Our model achieves competitive image understanding with far fewer visual tokens, making it better suited for multi-image scenarios.
-Fine-tuning Results of Referring Expression Comprehension.
-Our model demonstrates state-of-the-art performance on various referring expression comprehension (REC) benchmarks, as shown in Tab.
-2.
-Compared to recent REC models like VisionLLM [90] and Shikra [10], our model outperforms on most datasets.
-Remarkably, even though we only use public REC data [42, 58] for fine-tuning, our model matches QWen-VL [6] which utilizes a extra 22M in-house grounding dataset and trains at higher resolution (448 vs 224 pixels).
-This shows the effectiveness of synchronized fine-grained image features for enhancing the REC capability of our model.
-Fine-tuning Results of Segmentation-to-Image Translation.
+Tab.
+4 shows the results on referring expression comprehension (REC) benchmarks.
+Our method outperforms other methods.
+Even though we only use public REC data [42, 59] for fine-tuning, MM-Interleaved matches Qwen-VL [6] which utilizes an extra 22M in-house grounding dataset and trains at higher resolution (448 vs.
+224 pixels used by ours).
+This shows the effectiveness of synchronized fine-grained image features for enhancing the REC capability of our method.
+Tab.
+5(a) shows the performance on segmentation-to-image translation, which is a capability not shown in other multi-modal LLMs.
 We follow the protocol proposed in ControlNet [102] by generating images from ADE20K ground truth semantic masks and use OneFormer [38] for segmenting the generated images.
-Standard mIoU is reported as a metric for measuring the semantic and structural fidelity of the generated images.
-We show the results in Tab.
-3(a).
+Standard mIoU metric is used for measuring the semantic and structural fidelity of the generated images.
 MM-Interleaved clearly outperforms other baselines including ControlNet by a large margin.
-Compared to ControlNet, our model could leverage the better representations learned by our generative modeling from large-scale pre-training data, which benefits the image generation tasks.
-Moreover, thanks to MMFS, our model is capable of generating realistic images with pixel-level precise alignment from a semantic label map, where simply relying on coarse Perceiver Resampler features fails on this task, which we will show in the ablation studies.
-ADE20K (GT) [38] VQGAN [20] LDM [73]
+Compared to ControlNet, MM-Interleaved could leverage the better representations learned by our generative modeling from large-scale pre-training data, which benefits the image translation tasks.
+Moreover, thanks to MMFS, MM-Interleaved is capable of generating realistic images with pixel-level precise alignment from a semantic label map, where simply relying on coarse Perceiver Resampler features fails on this task (see Tab.
+6(c)).
+Groundtruth VQGAN [19] LDM [75]
 0.58 0.21 0.31
@@
 0.26 0.35 0.44
-(a) Segmentation-to-image generation results on ADE20K [106] mIoU ($\uparrow$) is reported.
+(a) Segmentation-to-image generation on ADE20K [108]. mIoU is reported.
 […]
@@
 MM-Interleaved 0.70 39.7
-(b) Image generation results on VIST [34].
+(b) Last image generation with interleaved context for visual storytelling on VIST [34].
 Model Pororo Flintstones
@@
 MM-Interleaved 14.7 18.7
-(c) Autoregressive image generation results on Pororo [49] and FlintStones [28]. FIDs($\downarrow$) are reported.
-Table 4: Comparative analysis of image generation ability of MM-Interleaved.
-Fine-tuning Results of Visual Storytelling.
-Besides text-to-image and image-to-image generation, our MM-Interleaved could also condition on interleaved image-text as context for generating new images.
-We evaluate our model’s performance both for generating the last image on the Visual StoryTelling (VIST) dataset [34] and for generating multiple images autoregressively on the Pororo [49] and Flintstones [28] datasets.
-As is shown in Tab.
-3(b) and  3(c), our model achieves SoTA performance on both CLIP-Similarity and FID metrics compared to previous works on VIST and achieves competitive performance compared with previous specialist methods on Pororo and Flinestones.
-More qualitative results and implementation details can be found in the appendix.
+(c) Multi-image generation with interleaved context for visual storytelling. FID is reported.
+Table 5: Segmentation-to-image generation and visual storytelling results.
+Tabs.
+5(b) and 5(c) show our results on visual storytelling, which is required to generate new images with interleaved image-text as context.
+We evaluate the performance both for generating the last image on the VIST dataset [34] and for generating multiple images auto-regressively on the Pororo [51] and Flintstones [28] datasets.
+Our model achieves SOTA performance even compared with previous specialist methods.
+Qualitative results can be found in Appendix.
+4.3 Ablation Study
+For ablation study, we use CLIP-ViT-L/14 [71] as the image encoder, OpenLLaMA-3B v2 [26] as the LLM and miniSD [^1] https://huggingface.co/justinpinkney/miniSD as the image decoder.
+The output image resolution is 256.
+The model is pre-trained for 10000 steps, using a mixture of LAION-COCO [78], LAION-2B [77], and MMC4 [111] datasets.
+For simplicity, MMFS in this subsection is by default implemented in a single-image and single-scale setting, where each token only attends to the $16\times 16$ feature map of the nearest preceding image.
+We evaluate the zero-shot performance on three representative tasks and four datasets, i.e., image captioning on the COCO Karpathy test set, text-to-image generation on the COCO Karpathy test set, visual question answering on OKVQA [60], and TextVQA [81] validation set.
+The evaluation metrics are CIDEr, FID-5k, and top-1 accuracy, respectively.
+To ablate the cross-attention module in image generation, we also evaluate the fine-tuned results for segmentation-to-image translation task on ADE20K [108].
 […]
@@
 256 110.9 29.5 29.6 27.8
-(a) pre-training with 224 input resolution
+(a) pre-training with 224 resolution
 […]
@@
 115.2 30.5 30.6 30.8
-(b) fine-tuning with 448 input resolution
+(b) fine-tuning with 448 resolution
 w/ MMFS ADE20$\uparrow$
@@
 (c) fine-tuning for image translation
-Table 5: Ablation on the usage of MMFS module.
-“Genetation” is the text-to-image generation task.
-“ADE20k” is the segmentation-to-image translation task.
+Table 6: Ablation on using MMFS.
+“Generation” is text-to-image generation task.
+“ADE20k” is segmentation-to-image translation task.
 Others are text generation tasks.
-“# Token” indicates the number of input visual tokens for LLM (32 by default).
+“# Token” indicates the number of input visual tokens for LLMs (32 by default).
 […]
@@
 Dense Resampler 32 tokens 107.2 30.7 28.9 24.0
-Table 6: Ablation on the design choice of MMFS module.
+Table 7: Ablation on the design choice of MMFS.
 Different attention modules can be used in MMFS.
 We also ablate whether to add additional transition layer before feeding image features into MMFS.
-Transformer attention with Resampler transition on single scale and single image is similar to the cross attention used in Flamingo.
-“Self-Attention” has the same parameters with “Resampler”.
-4.3 Ablation Study
-In the ablation study, we use CLIP-ViT-L/14 [70] as the vision foundation model, OpenLLaMA-3B v2 [26] as the multi-modal LLM and miniSD [^1] https://huggingface.co/justinpinkney/miniSD as the image decoder.
-The output resolution of each image is 256.
-For simplicity, in this subsection, we study the effect of token efficiency, image resolution, cross-attention mechanisms, and the role of MMFS in image generation in a single-image setting, where each token only attends to the nearest preceding image.
-Please refer to the appendix for more detailed ablations studies in the multi-image setting.
-The model is pre-trained for 10k steps on 80 NVIDIA 80G-A100 GPUs.
-The mini batch size of LAION-COCO [77], LAION-2B [76], and MMC4 [109] on a single GPU is 2, 2, and 4 respectively.
-For token efficiency, image resolution, and cross-attention mechanism, we evaluate the model’s zero-shot performance on three representative tasks and four datasets, i.e., image caption on the COCO Karpathy test set, text-to-image generation on the COCO Karpathy test set, visual question answering on OKVQA [59], and TextVQA validation set.
-We use CIDEr, FID-5k, and top-1 accuracy as the evaluation metrics.
-For cross-attention in image generation, we perform supervised fine-tuning on the ADE20K [106] for the label-to-image generation task.
+Dense cross-attention with Resampler transition on single scale and single image is similar to the cross attention used in Flamingo [3].
 Token Efficiency.
 As shown in Tab.
-4(a), our method achieves better overall performance while only using 32 visual tokens for each input image.
+6(a), when equipped with MMFS, using only 32 visual tokens per image can outperform using 256 visual tokens without MMFS.
 Such results demonstrate the effectiveness of our method when the context length is limited.
-To avoid the potential information bottleneck caused by the Resampler, we further increase the number of visual tokens to 256.
-Our method still outperforms the baseline in this setting.
-Scaling up Input Image Resolutions.
-The performance gap between our method and the baseline becomes larger when further increasing the input image resolution from 224 to 448, as shown in Tab.
-4(b).
-Such results indicate our method could better exploit the additional information gained from high resolution.
+As shown in Tab.
+6(b), the performance improvement of using MMFS becomes larger when increasing the input image resolution from 224 to 448.
+Such results indicate our method could better exploit the additional information gained from high resolution even with only 32 visual tokens.
+MMFS for Image Generation.
+Tab.
+6(c) demonstrates the criticality of adding MMFS for segmentation-to-image translation task.
+This task is hard as it requires precise pixel-level information to align the given segmentation condition and image output properly.
+Without using MMFS, this task fails, showing extremely low mIoU results.
+Visualizations results in Appendix shows that the generated image without MMFS cannot preserve all spatial information, and the spatial alignment for the generated results is poor.
 Comparison between Different Cross Attention Mechanisms.
 The ablation of adopting different cross-attention mechanisms for LLM is shown in Tab.
-6.
-The overall performance drops when directly replacing MMFS with the vanilla dense cross attention, possibly due to its slower convergence speed on 2D data, and using a resampler performs equally well.
+7.
+The overall performance drops when directly replacing MMFS with the vanilla dense cross-attention, possibly due to its slower convergence speed.
 We highlight that the model with Deformable attention performs significantly better than other attention mechanisms on TextVQA, indicating that Deformable attention could effectively and efficiently capture fine-grained information like text needed for the task, such as visual question answering in this case.
-MMFS for Image Generation.
-We study whether adding MMFS is critical for image generation on the segmentation-to-image task.
-This task is hard as it requires precious pixel-level information to align the given segmentation condition and image output properly.
-From the results in Tab.
-4(c), it is clear that the model without MMFS failed to generate images spatially aligned with the segmentation image condition, showing extremely low mIoU results.
-We show visualizations of the generated results in the supplementary material to provide a visual comparison.
-It is worth noting that the quality of the generated image from the model without MMFS is still high and properly aligned with the text description input; however, without MMFS, the model cannot preserve all spatial information, and the spatial alignment for the generated results is poor.
-5 Conclusion
-In conclusion, this paper presented MM-Interleaved, an end-to-end generative model specifically designed for interleaved image-text data.
-The unique contribution of MM-Interleaved lies in the novel multi-scale and multi-image feature synchronizer module, which significantly enhances the model’s capability to access and integrate fine-grained image features.
-This design addresses the longstanding limitation in existing models where a fixed number of visual tokens inadequately capture intricate image details, especially in scenarios involving multiple images.
-Pre-trained on diverse image-text datasets and further refined through supervised fine-tuning, MM-Interleaved shows a respectful proficiency in processing complex multi-modal instructions.
-The empirical results modestly highlight its ability to discern intricate visual details and generate images that align accurately with both textual and visual inputs.
-Appendix A Additional Ablation Studies
-This section provides more ablation studies for MM-Interleaved, all of which share the same settings as those in the main text by default.
-Note that these models are not fine-tuned on downstream tasks.
-Pre-training with Different Loss Terms.
-As showed in Tab.
-7, MM-Interleaved achieves better performance when jointly trained with both $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ on comprehension and generation tasks, demonstrating the mutual benefits between the two loss terms.
-Moreover, setting $\lambda=10$ achieves a better balance between $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ empirically.
-[…]
-$\mathcal{L}_{NTP}+100~{}\mathcal{L}_{NIP}$ 106.2 31.1 29.8 24.5
-$\mathcal{L}_{NTP}+10~{}\mathcal{L}_{NIP}$ 110.6 30.0 29.8 27.7
-$\mathcal{L}_{NTP}+\mathcal{L}_{NIP}$ 110.0 31.4 29.3 26.0
-$\mathcal{L}_{NTP}$ only 105.7 29.9 27.6
-$\mathcal{L}_{NIP}$ only 34.2
-Table 7: Pre-training with different loss terms.
-The Relationship between MMFS and resampler.
-The results in Tab.
-8 validate the mutual benefits between our proposed MMFS module and the Resampler used in image tokenizer, as removing either of them leads to an overall performance degradation.
-[…]
-110.6 30.0 29.8 27.7
-107.0 32.2 28.7 22.5
-102.7 32.0 27.3 22.0
-Table 8: The complementary relationship between MMFS and Resampler. When not using Resampler, we directly feed 32 randomly-initialized learnable embeddings as input visual tokens into the LLM.
 MMFS with Multi-Image and Multi-Scale.
 As shown in Tab.
-9, adding multi-image and multi-scale for MMFS improves the model’s performance on VQA tasks.
-To better take advantage of the multi-image mechanism, we follow [3, 83] to further evaluate our model with few-shot prompts.
-Two training samples including both image and text are selected as the input prompts.
-The results in Tab.
-10 demonstrates that MMFS with multi-image and multi-scale benefits the most from the few-shot setting, indicating its potential capability for interleaved image-text modeling.
+8, adding multi-image and multi-scale for MMFS improves the performance.
+To better take advantage of the multi-image mechanism, we further evaluate our model with few-shot prompts following [3, 84].
+As is illustrated in Fig.
+6 (left), when the number of context images increases, MMFS continuously outperforms only using 32 visual tokens, and also benefits from further attending to the multi-scale feature maps of multiple images.
 […]
@@
 115.8 30.0 31.7 32.0
-Table 9: Zero-shot evaluation results of MMFS.
-[…]
-36.5 30.7
-38.5 31.1
-38.7 33.1
-finetuning with 448 input resolution
-38.3 36.6
-38.7 37.7
-39.1 37.9
-Table 10: 2-shot evaluation results of MMFS on VQA tasks.
-Appendix B Implementation Details
-B.1 Pre-training
+Table 8: Ablation on multi-scale and multi-image usage of MMFS.
+Figure 6: Left: Few-shot results on OKVQA and TextVQA.
+Right: Additional GFLOPs over using 32 visual tokens with different numbers of image and text inputs.
+$N_{v}$ and $k$ denote the number of visual-token per image and the number of few-shot examples, respectively.
+$N_{i}$ and $N_{t}$ denote the number of image and the number of subsequent text tokens for each image in the sequence.
+Note that $N_{v}=256$ does not support $N_{i}\geq$ 8 images due to 2048 LLM context length.
+Computational Efficiency.
+Compared to only using 32 visual tokens, further integrating MMFS into LLM only increases by about 2%, 2%, 6% and 3% in FLOPs, #parameter, runtime and memory, respectively.
+When compared to using more visual tokens (i.e., 256) w/o MMFS, using MMFS with 32 visual tokens achieve better performance (see Tab.
+6(a)) and is much more efficient, with 2.8$\times$ fewer FLOPs and 1.3$\times$ fewer runtime.
+Fig.
+6 (right) shows the additional FLOPs over only using 32 visual tokens with synthesized interleaved image-text inputs.
+Our method consistently achieves lower or similar computation than using 256 visual tokens or using dense cross-attention.
+5 Conclusion
+This paper introduces MM-Interleaved, an end-to-end trained generative model for interleaved image-text comprehension and generation.
+It is built upon the Multi-Modal Feature Synchronizer (MMFS), which is proposed to enhance multi-modal LLMs by reducing the required visual tokens, enabling efficient extraction of visual details.
+Without using in-house data, our method achieves SOTA performance on various multi-modal benchmarks.
+We hope this work can contribute to the research on multi-modal LLMs for interleaved image-text scenarios.
+Limitations. Both the quality and quantity of public interleaved image-text data are relatively low, making it difficult to exploit the full potential of interleaved generative modeling.
+Negative impacts.
+As other multi-modal models, it may suffer from hallucination issues and potentially generate biased contents due to the noisy training data.
+We hope to invest more efforts to improve the quality and quantity of interleaved image-text data, with the aim of further improvements while ensuring its safety and reliability.
+Appendix 0.A Implementation Details
+0.A.1 Pre-training
 Dataset Details.
@@
 Data Concatenation Strategy.
-For image-text-pair datasets (e.g., LAION-2B, LAION-COCO [77]), we randomly sample multiple image-text pairs from the same dataset and concatenate them to the maximum context length (i.e., 2048) during pre-training.
+For image-text-pair datasets (i.e., LAION-2B, LAION-COCO [78]), we randomly sample multiple image-text pairs from the same dataset and concatenate them to the maximum context length (i.e., 2048) during pre-training.
 For interleaved image and text datasets (i.e., MMC4 [111]), we also split and concatenate the documents to form the training samples.
@@
 The detailed hyper-parameters of pre-training are listed in Tab.
-11.
-Besides that, for image generation, we ignore the training loss of images which are the first element in the sequence.
+9.
+Besides that, for image generation, we ignore the training loss of images that are the first element in the sequence.
 The text condition of the rest images are dropped with a 10% probability to improve classifier-free guidance sampling [32].
@@
 Warmup steps 1k
-Learning rate schedule cosine with warmup
+Learning rate schedule constant with warmup
 Training iterations 15k
@@
 Augmentation CenterCrop
-Batch size (per GPU) 2,2,4
+Batch size (per GPU) 2 for MMC4, and 4 for the rest
 Table 9: Hyper-parameters for pre-training.
-B.2 Supervised Fine-tuning
+0.A.2 Supervised Fine-tuning
 VQA and Image Captioning.
@@
 Referring Expression Comprehension.
-Following previous works [10, 6], we train our model in the form of VQA, i.e., the prompt being Question: Provide the bounding box coordinate of the region this sentence describes: {object}, Answer: (x1,y1)(x2,y2).
+Following previous works [10, 6], we train our model in the form of question answering, i.e., the prompt being Question: Provide the bounding box coordinate of the region this sentence
+describes: {object}, Answer: (x1,y1)(x2,y2).
 The generated bounding box is considered correct if its intersection over union (IoU) with the GT box is greater than 0.5.
@@
 Table 12: Hyper-parameters for Segmentation-to-Image Translation.
-Visual Storytelling. Following previous works [43, 104, 66], MM-Interleaved is finetuned on the following three visual storytelling datasets respectively:
+Visual Storytelling.
+Following previous works [43, 106, 67], MM-Interleaved is finetuned on the following three visual storytelling datasets respectively.
+The finetuning hyper-parameters are listed in Tab.
+13.
   • VIST [34] is a real-world vision-language dataset, containing 34k and 5k samples for training and evaluation.
@@
 FID [31] is used as the evaluation metric.
-The finetuning hyper-parameters are listed in Tab. 15.
 Hyper-parameters VIST / PororoSV / FlintstonesSV
@@
 Table 13: Hyper-parameters for visual storytelling.
-B.3 Evaluation
+0.A.3 Evaluation
 Dataset Task Split Metric
@@
 Generation. MS-COCO [52] Text-to-image generation val-30K FID($\downarrow$[31]
-LN-COCO [69] Text-to-image generation val-30k FID($\downarrow$[31]
+LN-COCO [70] Text-to-image generation val FID($\downarrow$[31]
 ADE20k [108] Segmentation-to-image generation val mIoU($\uparrow$)
@@
 All these evaluation tasks and metrics are listed in Tab.
-16.
+14.
 Image Generation. For all image generation tasks, the scale of classifier-free guidance [32] and the total inference step is set as 3.5 and 250 by default.
@@
 The prompt templates for each text generation tasks are listed in Tab.
-17.
+15.
 The ‘Image Caption (short)’ task includes COCO Caption [11], Flicker30k [69] and NoCaps [2], while the ‘Image Caption (long)’ task includes Image2Paragraph [44].
-Task
-Prompt Template
-Zero-shot
-Image Caption (short)
-{image} a photo of
-Image Caption (long)
-{image} Please describe the image in detail. The image depicts
-VQA (except VizWiz)
-Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
-VizWiz QA
-Based on the image, please answer the question.
+Task Prompt Template
+Zero-shot Image Caption (short) {image} a photo of
+Image Caption (long) {image} Please describe the image in detail. The image depicts
+VQA (except VizWiz) Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
+VizWiz QA Based on the image, please answer the question.
 {image}{question} When the provided information is insufficient, respond with ’Unanswerable’.
@@
 The answer is:
-Visual Dialog
 […]
-Supervised Fine-tuning
-Image Caption (short)
-{image} Provide a one-sentence caption for the provided image.
-Image Caption (long)
-{image} Please describe the image in detail.
-VQA (except VizWiz)
-Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
-VizWiz VQA
-Based on the image, please answer the question.
+Supervised Fine-tuning Image Caption (short) {image} Provide a one-sentence caption for the provided image.
+Image Caption (long) {image} Please describe the image in detail.
+VQA (except VizWiz) Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
+VizWiz VQA Based on the image, please answer the question.
 {image}{question} When the provided information is insufficient, respond with ’Unanswerable’.
@@
 The answer is:
-Visual Dialog
 […]
 Table 15: Prompt templates for text generation.
-Appendix C Qualitative Results
-C.1 VL Comprehension
+Appendix 0.B Additional Ablation Studies
+This section provides more ablation studies for MM-Interleaved, all of which share the same settings as those in the main text by default.
+Note that these models are not fine-tuned on downstream tasks.
+Pre-training with Different Loss Terms.
+As shown in Tab.
+16, MM-Interleaved achieves better performance when jointly trained with both $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ on comprehension and generation tasks, demonstrating the mutual benefits between the two loss terms.
+Moreover, setting $\lambda=10$ achieves a better balance between $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ empirically.
+[…]
+$\mathcal{L}_{NTP}+100~{}\mathcal{L}_{NIP}$ 106.2 31.1 29.8 24.5
+$\mathcal{L}_{NTP}+10~{}\mathcal{L}_{NIP}$ 110.6 30.0 29.8 27.7
+$\mathcal{L}_{NTP}+\mathcal{L}_{NIP}$ 110.0 31.4 29.3 26.0
+$\mathcal{L}_{NTP}$ only 105.7 29.9 27.6
+$\mathcal{L}_{NIP}$ only 34.2
+Table 16: Pre-training with different loss terms.
+The Relationship between MMFS and Resampler.
+The results in Tab.
+17 validate the mutual benefits between our proposed MMFS module and the Resampler used in image tokenizer, as removing either of them leads to an overall performance degradation.
+[…]
+110.6 30.0 29.8 27.7
+107.0 32.2 28.7 22.5
+102.7 32.0 27.3 22.0
+Table 17: The complementary relationship between MMFS and Resampler. When not using Resampler, we directly feed 32 randomly-initialized learnable embeddings as input visual tokens into the LLM.
+Appendix 0.C Qualitative Results
+0.C.1 Zero-shot Results
+The zero-shot comprehension and generation visualizations of our method are shown in Figs.
+7, 8, 9 and 10.
+Specifically, Fig.
+7 demonstrates the emergent multi-modal in-context learning capability of our method, while Fig.
+8 further illustrates the effectiveness of our method in understanding complex scenarios such as robotics, gaming and GUI interface.
+Fig.
+9 exhibits that our method can also generate appropriate images based on the provided contexts of desired styles or concepts.
+Fig.
+10 displays our method’s capability of generating images and texts interleavedly.
+Figure 7: Zero-shot text generation with interleaved images and texts serving as multi-modal few-shot contexts.
+Figure 8: Zero-shot text generation with interleaved images and texts on complex scenarios such as robotics, gaming and GUI interface.
+Figure 9: Zero-shot image generation with interleaved images and texts. Our method can generate appropriate images based on the provided contexts of desired styles or concepts.
+Figure 10: Zero-shot interleaved image and text generation for visual storytelling and multi-modal instructions.
+0.C.2 Supervised Finetuning Results
 Text Reading QA. As is shown in Fig. 11, MM-Interleaved with MMFS provides more accurate answers when requiring fine-grained details for generating text outputs given VL inputs.
@@
 The visualization of MM-Interleaved on REC tasks is shown in Fig.
-6.
+12.
 Our model with MMFS is also capable of generating more accurate coordinates given the referring expression and the query image.
@@
 The image shapes are normalized for visualization.
-C.2 Image Generation
 Segmentation-to-image Translation.
@@
 In Fig.
-8, we compare the multiple images sequentially generated by MM-Interleaved with and without MMFS.
+14, we compare the multiple images sequentially generated by MM-Interleaved with and without MMFS.
 The images generated with MMFS achieves better spatial consistency (e.g.
@@
 As is shown in Fig.
-9, given the first frame image and caption as context, MM-Interleaved with MMFS generates the following interleaved images and texts coherently, achieving balance between the generation diversity and spatial semantic consistency.
+15, given the first frame image and caption as context, MM-Interleaved with MMFS generates the following interleaved images and texts coherently, achieving balance between the generation diversity and spatial semantic consistency.
 Figure 14: Multi-image Generation on PororoSV [51] and FlintstonesSV [28].
@@
 Only the caption and image of the first frame is given as condition during generation.
-Appendix D Future Work
-Future works of MM-Interleaved aims to further scale up both the model and data size, enabling end-to-end training with full-parameters from scratch.
-In addition, we believe that it is also worth exploring to build a comprehensive evaluation benchmark specifically designed for interleaved image-text modeling.